A Complete Bibliography of Publications in the European Journal for Philosophy of Science

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

27 May 2024 Version 1.00

Title word cross-reference

H → γγ [Pas19]. N [SM23].

*the* [LC21].

-of-1 [SM23].

19 [Bir21, Bir23, DL22, LC21].

5 [GJB23].

abduction [FEG19]. abstract [CK23]. Abstraction [Rae18, FM18, Zac22]. academic [KM16a, Kro21]. accelerating [Gur20]. accessing [CH23]. accident [Wan24]. account [Cla13, Dim20a, Dow19, HS19b, Kro21, Mor21, Reu20, Spo14, Tun22, Wan24]. accounts [Fri22, Moh22]. across [KM16b, IW12]. Action [Dow19, MM16]. active [Bec12]. activist [Kos22]. activities [Joh12]. activity [Aus20]. adaptation [GLB15, Hea15]. Adapting [GAD^{+}21]. Adding [BDPP15, Bou24]. Addressing [IdMM14]. adequacy [BC17, BP21b, LE22]. adequacy-for-purpose [BP21b, LE22]. admissibility [LWW21]. advisory [Lew19]. aesthetic [Mon13]. Afactivism [Tay23b]. affected [Reu20]. afraid [FV11]. after [HM19, GS16]. again [Raa21]. Against [All23, Lau14, Laz18, Par19, Rov17]. agent [Dou23]. agent-based [Dou23]. agential [Jak23a]. agreement [Det24]. AI [Pan24]. algorithmically [SR22]. Aligning [BP21a]. alignment [Can22]. alternative [Kos17, Trl22]. alternatives [Her14, Kas23]. Aluminium [Hij22]. always [FV11, GS13]. Amending [Bor11]. analogies [McC20, Nap21]. Analogy [Hij22]. Analysis [FS21, BCVV21, BC19, Han16, Lis17, LT21, dRK21]. analytic [FM18]. analytical [AJ17, Sol11a]. ancient [VG22]. animal [VB21]. Anthropocene [LR24]. anthropogenic [GPM23]. anthropological [Kid21]. Anti [Ros13]. Anti-Darwinism [Ros13]. antidote [FS22, NNW19]. apology [Jak21a]. Aporia [Sal17]. appealing [Hac22]. Appearance [Gio23, Jak23a]. applicability [Mac21b, PUB22]. application [VH21]. application-oriented [VH21]. applications [Mol22, Mus22]. Applied [VG22, PES22]. apply [Reu17]. Applying [EW22, SW21, SOH21, PES22]. appraisal [Dun22]. appraised [Sol17]. approach [Can22, De 23, DS23, Har19, OÖ21, Val14]. approaches [Bas21, Kha22]. architectures [SB15]. archival [BS21]. Argument [Bac13, Ala12, Bri14, Ero12, Hen21, Her14, Par21, Rov17, Spr16]. argumentation [EW22, IN22]. arguments [Han22, Jak23b, Ney23, Par19, Pet16, Sch13]. Armstrongian [Moh22]. art [Ans21]. artifacts [Fee22, Knu21]. artifactualism [dO22]. Artificial [Lut12, Anz21, Mor21, Pan24]. arts [VN21]. ask [FV11]. aspects [Ben21, BC22, Sol11a]. Assessing [KP15]. assessment [GPM23, LE22]. association [Spe18]. astrophysical [CS23]. asymptotic [Pex14]. atom [BP15]. atoms [CH23]. Attention [Tay23a]. attitude [Gra24]. attitudes [Spe18]. attribution [GPM23]. audience [Kit19]. automaton [FD21]. Autonoesis [And23]. autonomy [Mei19, Sch15]. Autopoiesis [Mei19]. autopsy [RW11]. Averaged [SM23]. avoid [DL22, Ros13]. avoidance [Lan19]. axiomatization [Odd24].

Bachelard [Sim18]. back [Men20, Sim18]. backgrounding [Hoc22]. backward [FM18, Lóp19]. backward-error [FM18]. Backwards [Fry16]. bad [Mus22]. Barad [EG22b]. bargaining [Fre23]. based [Bas21, BCVV21, DG23, Dou23, GAD^{+}21, LT21, Sol11b, VDE17]. basis [GY22]. Battle [Mar19, Mar20]. Bayes [Mas21a]. Bayesian [Bar18, CC23, FEG19, Lae11, Spr13]. Bayesianism [Fis21, Lae11, Spr18, WW20]. be [Acu21, BL15, BC19, GJB23, GS13, Pex14, SH19]. Beauty [Mon13, Koz23]. becoming [Aam22]. beginning [Lin21]. behavioral [BG21]. being [Aus20]. belief [Luk21]. Bell [GH23, Lau18]. benefits [Pet21]. Benovsky [Maz16]. Best [Dop14, Díe18, Dow19, Dra17, Fri22, Sag23, Sch15]. Better [BS21]. between [Bas21, Dim20a, GL21, Int15, Jak21a, Joh12, KL16, KL23, KG14, LL18, Luc23, McC20, VG22]. Betz [Joh15a, Joh15b]. beyond [Adl22, Dup11, Hil22, Koz23, Pes11, WR23]. biannual [RTT19]. bias [MK21, Reu20]. Big [LR20]. binary [Kar21]. Biodiversity [Boc22]. Biological [Art21, Au21, IK21, JFEH19, Kos19, Mei19]. biology [BDPP15, BC22, BBŠ21, IK21, JFEH19, Kos17, MN16, Mac21b, Sch13, Zac22]. Bipedal [Bak21]. birds [Shr14]. black [HM21, RP14, Sen17]. blue [Wal23]. blurry [RU18a, RU18b]. Bohmian [Sol17]. Bohr [BP15]. books [WG17]. boolean [Har19]. bouncing [ET20]. boundaries [LR24, WR23]. brakes [Fis24]. breaking [Fri13]. bridging [VG22]. bringing [Rey21]. Broken [Fis24]. build [NP18, Nor18a, Nor18b].

Call [AF22]. Calling [BS20]. Can [Maz14, Pex14, Bei12, BL15, BC19, ET20, IdMM14, Ros13, Şer17]. cancer [BC19, Plu21]. candidate [Kin20]. capacities [Joh12]. Capturing [Anz21]. Carnap [Jus12, Sar13, Ueb11]. Carnot [KB11]. Cartwright [Kri20, Kri21, Näg22]. Case [Hij22, Bak21, BS20, BvE21, De 21, Fee20, Fri22, GS13, Han16, Hea15, Kar21, KB11, KM16b, Lal23, Lew19, Mät22, RW11, Sha21, SOH21, Sul19]. catch [Sch15]. catch-22 [Sch15]. categoricity [Toa21]. Category [Eva16]. Causal [Dim20b, Aam22, Bas21, Bou19, BBŠ21, Ero12, Fis24, GS13, GL21, Hen18, Kro21, Kuo12, LOP18, Reu17, SR13, Shr14, Sob20, VJ19, Web22, Woo22, Fri12]. causality [Bou24, RW11]. Causation [San12, Dim20b, Web22]. cause [EE14]. caused [Kõi23, Sul19]. causes [Fri12, Näg22]. cell [Dup11]. cells [Suá23]. centered [Wan24]. century [Luc22, Hij22]. certainty [Sch22b]. ceteris [Spo14]. challenge [Bec12, BCVV21, Bog21]. challenges [Fum16, Har19]. Chang [El 21]. change [Can22, De 23, GPM23, MZG17]. changeable [RU18a, RU18b]. changes [Pen21]. Changing [SGH21]. chaos [NNW19]. checking [FRT24]. Chemistry [Hij22]. chimera [Yee21]. Chinese [Sha21]. Choice [Fum16, AD14, BFŠS19, Pet18, Wel11]. Chunk [BP15]. Church [De 21]. citizen [Yee23]. civil [Mac21c]. clarification [Fri13, Spe18]. clarifying [Bar22]. clash [OC11]. classes [Det24]. Classicality [GH23]. Classification [RE22, Bas21, Kar21, Lud23]. classificatory [Bas21]. Classifying [Mät22]. clear [Str21]. Climate [BL15, Lam21, Bet15, Det24, Fri15, GPM23, Han20, Hea15, Int15, JB21a, KP15, Kat23, O’L23, PUB22, SS15, VNT22]. clock [Val16]. Close [Nap21]. closure [Dim20b]. cluster [RU18a, RU18b]. coarse [Mor23]. cognition [BM23, Fee20, Tay23b]. cognitive [Bed19, Ben21, CC23, MK21, Wel11]. coherence [El 21]. Coherent [Web22, SD21]. coherentism [Mor19]. coin [Mol22, Par21]. coincidence [Gio21, Gio22]. coincidences [Gio21, Gio22]. coins [How17]. collection [Mas21b, Zah23]. collective [Ans21, Hue11, Pol24]. comments [RE24]. common [EE14, Näg22]. communication [GJB23, Hen20]. community [Pol21]. comparative [She21, dRK21]. compensatory [Aiz13]. complementary [Rae18]. Complex [Suá22, LLW13, Plu21, SS15]. complexities [dMMI23]. Complexity [BC19]. component [Lus22]. components [PG20]. Composition [Hij22]. compulsions [Sar13]. computational [HS19a, HW21, Mac21b]. computer [Bei12, Bei18, Bog19, Bog22]. conceivability [All22]. concept [Jus12, SD20, She21, Tay23a]. conception [Wri12]. conceptions [Dup11, IdMM14]. Concepts [AF22, Kor18, SFG^{+}18a, SFG^{+}18b]. Conceptual [Spe18, TV17, De 23, MZG17, OÖ21, Rey21]. concerning [Sob20]. condition [MP23, Roc12, Roc17]. conditionals [Fri16]. conditions [Hei14, Spo14]. conference [RTT19]. confidence [Mac21a]. confidently [Kho23]. Confirmation [Roc16, Lae11, Sch22b]. conflicts [Jak21a]. conjecture [Wro14]. conjunction [Sch22b]. consciousness [All22]. consequences [KK21, Sol11a]. considerations [GJB23]. consistency [CDR23, Gur20]. conspiracy [San12]. constitution [Pov20]. Constitutive [San21, Bae12, Har19]. constraining [Fel21]. constraint [Sch22a]. Construct [Fee20]. constructionism [Han20]. constructions [FX21]. context [Auf11, Bar18, Kat23, Sol11b]. Contextual [Bor11, Acu21]. Contextualist [VFHS23]. continuity [All23, MZG17]. continuous [Maz14, Wer11]. continuous-time [Wer11]. continuum [Trl22]. contraction [Gio23]. Contrast [Det24]. Contrastive [Kuo12, Sob20]. contributions [VNT22]. control [Web22]. controversies [Auf11]. convergence [Sch22b]. Convergent [Pow12]. converse [Mol22]. cooperation [JP21]. coordinates [Gio21, Gio22]. Coordination [Mül24]. Coproduction [AF22]. Correction [Bir23, Bog22, GS22a, Gio22, KL23, Kri21, Mar20, NP18, RU18a, SFG^{+}18a, Sto21a]. Cosmic [Mas21a, Rya24, Sim21]. cosmology [DM23]. costs [Pet21, Bak21]. Countability [Hea21]. counterfactual [Reu17, Sch22a]. counterfactuals [Lan19]. counts [Fle12]. course [Mac21c]. courses [Lus22, You22]. COVID [Bir21, Bir23, DL22, LC21]. COVID-19 [Bir21, Bir23, DL22, LC21]. craniology [Luc22]. creative [FEG19, She21]. Creativity [Chu22, FX21, Kid21, Ans21, Jak21b, Kro21, Mor21, SD20, She21]. credible [Bet15, Han16]. crises [Sal17]. crisis [Hen20]. Criteria [Wan24, CH23]. critical [Fri15, GPM23, Bor11]. criticism [Sta12]. criticize [Zac22]. Critique [CC23, Hei14, Rub20, Suá14]. crucial [And23]. Cultural [VNT22, Han22]. current [Bsc20]. curriculum [Lus22]. curse [GY22].

D [NP18]. dangerous [DL22]. dark [Mas21a]. Darwin [Hel17, PS18]. Darwinism [Ros13]. Data [BP21b, Fee22, And23, Ant21, Leo19, Mac21a, Zah23, LR20]. Datasets [Mas21a]. Dawid [Her14]. day [SH19]. day-to-day [SH19]. Dealing [RU18a, RU18b]. debate [Bed19, GJB23, Gio23, Luc23, Näg22]. Decision [Wel11, Hea15]. decision-making [Hea15]. Decoherence [Jak23a].

Decoupling [Cro15, For24]. deduction [Hub17]. defence [Bet13]. defended [Kha15, Ueb11]. Defending [Le 12, Bor11]. Defense [FS21, Bel19, Moh22]. defining [Hac22]. Definition [Cel18]. definitions [Dob19]. Deflating [LFL16]. deflationary [LFL16]. degenerative [Dun22]. Democratising [AF22]. demonstration [CH23]. denial [Han20]. depend [VJ19]. Dependence [Vas20, Fag22, Kro11]. depression [Run21]. Derivational [Han16]. descriptions [Wer11]. Design [GLB15, Fle19, Kos19, Mac21c, SM23]. details [Bae12]. detecting [Pet21]. determination [Jus12]. Determinism [Adl22, Dob19]. deterministic [Sob20, Wer11]. detrimental [BL15, Pae24]. devo [Aus17]. diachronic [Big20, Sar15]. diagnoses [Fel21]. Diagnosing [O’L23]. diagrams [Pas19]. dialogue [Dan21]. diamonds [Aam22]. dichotomy [PES22]. did [Pas19]. difference [PW18, Sch15]. differences [Aiz13]. different [GS16, JB21b, MM18]. differentiated [Dup11]. Diffracting [EG22b]. dilemma [SB15]. dimension [Bri14, Hen20, Knu21]. dimensions [Bed19, Kro11]. Direct [Mol22, Str21]. direction [RR24]. disciplinary [Boo20]. disciplines [KL16, KL23]. discovery [Kat23]. discrete [Maz14]. discrimination [Roc16]. disease [Dup11]. disorder [GJB23]. dispositional [Hir21, Moh22]. dispositions [Aus17, Suá23]. dissent [BL15, Pae24]. dissipative [OÖ21]. distance [Dow19]. distant [Vez17]. distinction [Auf11, HLMV21, Mor23, Web22]. distinctively [Pov20]. distinguishes [Leo19]. Distinguishing [Int15, Jak23b, LL18]. distribution [Hel17]. diversity [BFŠS19, Pet21, SFG^{+}18a, SFG^{+}18b]. division [Bed19, Lew19]. Do [Pet16, DS23, GM22, KB11]. Does [Lae11, OC11, Reu17, Joh15a, Joh15b, VJ19]. doesn’t [Häg22]. doing [McL11]. dome [Fle12]. donation [Lew19]. Double [Hen20]. doubt [BL15, SH19]. Down [Kle21, Ros13]. dreaming [Fis24]. driven [IdMM14]. drivers [Fis24]. droplets [ET20]. DSM [GJB23]. DSM-5 [GJB23]. Dutch [WG17]. dynamic [TB21]. dynamical [AJ21, HLMV21]. dynamics [Sal17, WR23].

each [Bei18]. early [Gio23, Hij22]. Earth [Kle21, LR24]. ecology [DS23, Rae18]. economic [Han16, KL16, KL23]. economics [VFHS23]. economists [Mar13]. economy [BP21a, PS18]. Econophysics [Yee21]. edges [RU18a, RU18b]. Editorial [DH11]. Editors [IR17, IR18a, IR18b]. education [Kle21, Lus22]. educational [BOS22]. effect [JB21b]. effective [Bai13, Saa22]. effects [Due23, GPM23, GM22]. Efficiency [Fre23, BFŠS19, SR22]. Egypt [VG22]. Einstein [Gio18, Gio21, Gio22, Gio23, Nor11]. elementary [Big20]. elements [San21]. elephant [HW21]. elimination [Lan20]. eliminativism [TV17]. eliminativist [Cla13]. emancipatory [Kos22]. Emergence [Bai13, Pex16, Bel22, BB15, Cro15, De 19, For24, Gho23, GS16, Sar15, Sei20]. Emergent [EG22b]. Emerging [Dup11, MN16]. Émile [Gio18]. emotions [Hue11]. empirical [AD14, BC17, Gom21, Lus22]. empiricism [Pet18, Sar13, Bor11]. empiricist [Suá14]. EMU [Kha15, New14]. encapsulation [SB15]. encounters [Nap21]. encroachment [Mil14]. endogenously [Bec12]. endorsement [Gur20]. energy [BP21a, Mas21a, SR22]. engagement [IN22]. engine [KB11]. engineering [BvE21, Bar22, Kos19, Mac21c, Wan24, vE15]. entanglement [CCS22]. entities [For24]. entropy [She20]. environment [HS19a]. epigenetics [Bsc20]. episode [Bol21]. Epistemic [Au21, HW21, Kat23, Knu21, RE22, EW22, FRT24, Kro11, KK21, LT21, LE22, Pet21, SD20, Spo14, SFG^{+}18a, SFG^{+}18b]. epistemically [BL15, Pae24, Reu20]. Epistemological [BOS22, THP23, Mas21b, Sol11a, Sol11b]. Epistemology [BV19, LOP18, Boo20, Cla13, DL22, Luc22, Pot20, You22]. EPR [Bac13, EE14, San12]. EPSA17 [RTT19]. equilibrium [Han16, Val14]. equivalence [AD14, Luc23, OC11, Wer11, WR23]. era [BC19, DL22]. Erratum [Joh15a, Nor18a]. error [FM18]. errors [O’L23]. essence [Häg22, HM19]. essentialism [Hir21]. eternalism [Le 20]. ethical [SFG^{+}18a, SFG^{+}18b]. ethics [BS21, Dan21]. Ethnobiological [RE24]. Euclidean [De 23]. Euler [Räz18]. evaluation [VFHS23]. evaluations [Mon13]. Everett [Bar11]. Everettian [GL24]. Everything [FV11]. Evidence [Gur20, LT21, Näg22, Au21, Aut16, BCVV21, Dra17, Fri15, Hen18, Lan20, PW18, Sag23, Sol11b, Spe18, Sta12, Sto21a, Sto21b, Vez17, Wal23]. evidence-based [BCVV21, Sol11b]. evidential [Reu20, SW21]. Evo [Aus17]. Evo-devo [Aus17]. evolution [Adl22, Mül24, PS18, Pow12, Rey21]. Evolutionary [GS22a, GS22b, Suá22, Con18, Sch13]. example [Joh15a, Joh15b]. exclusion [Ero12, Woo22]. Exeter [RTT19]. exhibit [MZG17]. expanding [Mül24]. experience [Dar22]. Experimental [CH23, Anz21, Fee22, Fle19, Hen20, Mac21b, Sol11a, Sto21a, Sto21b]. experimentation [Mät22, Pot20, Sto21a, Sto21b]. Experimenter [FD21]. experiments [Bae12, Bei18, Bog19, Bog22, Cla13, Han19, Hit12, JB21b, Kho23, Koz23, Pas19]. expertise [Hit12, Tun22]. explain [DS23, Mar13]. Explaining [Bar19, Far20, Van14]. Explanation [And23, FM18, Pot11, Bak21, Ban17, BS20, BvE21, BC22, CPW20, Dra17, Fry16, Kri20, Kri21, Kuo12, Pov21, Reu17, Sag23, Sch22a, Sch15, Şer17, Shr14, Sob20, Sul19, Van14, VJ19, Vir21, Wri12, Zac22, vE15]. Explanations [Kin20, Bar22, BM23, DS23, EE14, HS19a, JFEH19, PG20, Pov20, Reu17]. explanatoriness [Sob20]. Explanatory [BBŠ21, Way18, Bar22, KG14, New14, Pex14, Räz18]. Explication [De 21]. Exploratory [Jac22, Mät22]. Exploring [IK21, dMMI23, FD21]. extended [Sch13, Tun22, GS22a, GS22b]. Extra [KM16a]. Extra-academic [KM16a]. Extrapolating [Kho23]. extremis [Bir21, Bir23]. extrinsicness [Suá23].

F [Pen21]. faced [BC19]. facets [Lop21]. fact [FRT24, Hel17]. fact-checking [FRT24]. facts [CDR23]. fairness [Fre23]. Farmer [Han19]. Fatal [Díe18]. fate [Ros13]. feminist [Auf11, IdMM14]. fertility [Sch17].

Feyerabend [Mar20, Mar19, Sha21]. Feyerabendian [Tam14]. Feynman [Pas19]. Fictional [Bar22]. field [Bai13, EWVTL24, For24, HR18, KM16b, Lam15, OÖ21, Sal17]. fields [Laz18]. figures [Pet21]. Filling [Bae12]. financial [VFHS23]. Fine [Bar18, Mor23]. fine-grained [Mor23]. Fine-tuning [Bar18]. Finetti [Ber19]. Fisher [Rub20]. Fitness [Suá22, TB21]. flexible [Dou23]. flip [Par21]. fly [Shr14]. Fodor [Ros13]. foes [Maz16]. Follow [LC21]. force [Bar19, Sha18]. Foregrounding [Hoc22]. forks [GS13]. form [Lar22]. Formal [Pol21]. formalism [LL18, tV22]. forms [DL22]. foundations [OÖ21]. fragility [Nor22]. fragmentation [TV17]. Frames [Kor18]. framework [VH21]. frameworks [Kob23]. Francis [Dan21]. Free [Joh15a, Joh15b, BP21a, Bet13, HH23]. free-energy [BP21a]. freedom [Kro21, MS23, Sha21]. Frege [GY22]. frontiers [Pan24]. fruitfulness [Iva19]. function [HR18, Ney23]. functional [DiF18, Gua23]. functionalism [BG21]. functions [Art21]. funding [Sha21]. funds [VFHS23]. future [HS12]. futures [Vir23].

Gait [Bak21]. Galilean [And23]. games [Fre23]. gap [VG22]. gases [KL16, KL23]. Gauge [Fri13, Teh15, BB15]. gene [Bel22]. General [Kas23, EG22a, Han16, HS19b, Vas20]. Generalized [Kob23]. Generic [RW11]. genetic [Kos17]. Genetically [Kõi23]. genuine [Sch22b]. Genuinely [Hue11]. geographical [Hel17]. Geometric [Bri14]. Geometry [Dar22, Luk21]. geoscience [Kle21]. Ghins [Ala12]. Giere [Bak20]. give [KB11]. Gleason [Acu21]. go [Lau14, Lóp19]. Gödel [Aam22]. good [Mus22, Zah23]. governing [ILP21]. grained [Mor23]. graining [Mor23]. grant [SR22]. gravity [Le 20, Pin15]. great [Hel17]. Gregor [Joh15a, Joh15b]. grounding [RE24]. Group [For24, Sul19]. groups [Cha20]. growth [Dun22]. guide [Bae22, Lóp23].

Hansson [VNT22]. happiness [Ang13]. hard [All22]. harmful [GM22]. Hartmann [Her14, Pin15]. Hasok [El 21]. Hausman [Näg22]. hawkmoths [FS22, NNW19]. head [Sim18]. health [Sal17]. healthcare [BS21]. heat [KB11]. help [IdMM14]. Hempelian [DS23]. Henri [Dan21]. hermeneutical [Kha22]. heterogeneity [Kuo12, Str21]. heuristic [Fis24, GY22, Kos19]. heuristics [Rae18]. Hidden [Pet21, Acu21, HS21]. hierarchies [Pex16]. hierarchy [BBŠ21]. high [Roc17, SR22]. Historical [Vic12, BC22, Gra24, Gua23]. historico [Dar22]. Histories [LL18]. historiography [Vir21]. History [Kle21, Nor11, BS21, Dim20a, Fel21, Hac22]. Hoefer [Häg22]. holes [Bar22, HM21, RP14, Sen17]. Holism [Gom21, Rae18]. homomorphism [PS16]. Hooft [BB15]. horizon [HM21, Pol24]. how-possibly [Kos17]. hows [dRK21]. HPS [Lus22]. human [FD21, MZG24]. Hume [Bel19]. Humean [Fri22]. Humeanism [Mul18]. hunt [Mas21a]. hybrid [HLMV21]. hydrogen [BP15]. hydroxychloroquine [Lal23]. hylomorphism [Sim21]. hypotheses [Lan20, Sob20, Str21]. hypothesis [Val16].

IBE [BvE21]. ideal [Bet13, KB11, PUB22, Joh15a, Joh15b]. idealization [And23, Zac22]. Idealizations [FS21]. Identifying [BD13]. identity [Big20, Hea21]. if [Bei18]. ignores [New14]. II [BP15]. illegitimate [Int15]. Imagination [Fre20, PUB22]. Immovable [Sha18]. implications [OÖ21, Sob20, SFG^{+}18a, SFG^{+}18b]. implicit [Fee20, Spe18]. implicit-association [Spe18]. imprecise [Fis21]. improbable [FX21]. improving [Bas21]. in-principle [Hen18]. inchworm [Bol21]. included [GJB23]. Incommensurability [GS22b, Boc22, GS22a]. Incompatibility [Bog21]. inconclusiveness [Hen18]. inconsistent [Pov21]. Incorporate [RE22]. increase [Roc16]. indecomposability [Gho23]. independence [Con18, Kro21, San14, MP23, San14]. independently [Bas21]. indeterminacy [GL24, Jac22, Old24]. indeterminism [Lar22, Sha22]. indeterministic [DG23, Wer11]. index [VFHS23]. indifference [Wil18]. individual [SM23]. individuality [Bou24]. individualized [SM23]. individuation [DiF18]. Induction [GS18, Bel19, Bog21, Hub17, Nor11]. Inductive [Pae24, Con21, Hen21, HH23, Kar21]. ineffective [FS22]. Inference [Dra17, Sch15, Har19, KK21, LOP18, New14, Sag23]. inferences [Bog19, Bog22]. infinite [NP18, Nor18a, Nor18b]. infinitely [How17]. Infinity [Trl22, CV20, Rya24, Sha18, Sha22]. Inflating [Ben21]. inflationary [DM23]. influence [Can22]. informal [LWW21]. Information [Pov21, She20, Bou19, Bou24, Lau23, LFL16]. information-theoretic [Bou24, Lau23]. informational [SB15]. initial [Bir21, Bir23]. inquiry [BFŠS19, Chu22]. insights [Tam14]. Insolubility [Bac13]. instability [AJ21]. instrumental [Bak20]. instrumentalist [Ber19]. Integrating [Mac21c]. integration [BC19, GY16, Way18]. Integrative [Con18, MN16, Mac21c]. intelligence [Luc22, She21]. interacting [Pex16]. interactive [GM22, Kõi23, Näg22]. intercomparisons [O’L23]. Interdisciplinarity [MM16, EWVTL24, LR24, Mäk16]. Interdisciplinary [GY16, MN16, BV19, EWVTL24]. interest [Lus22]. interests [Hil22]. internal [Sal17]. interpretation [Esf13, Hoc22, LL18, Pas19, Sar13]. interpretationalism [Luc23]. Interpreting [DdR18]. Interpretive [McC20]. intervention [BD13]. interventions [Woo22]. Introduction [ACBB11, KP15, Mas21b, HS12]. investigate [SM23]. investigations [Wan24]. invisible [Pet21]. IPCC [Joh15a, Joh15b]. Irrational [Gho23]. Irresistible [Sha18]. irreversibility [tV22]. issues [BOS22, HW21].

Jan [Her14]. Janus [Sim18]. Jerry [Ros13]. John [NP18]. journals [Dun22]. Just [Sol11b]. justification [Hub17]. Justifying [Wil18].

kind [GJB23, Gio18, Kõi23, Nap21, Raa21]. kinds [Fag22, Gua23, RE24, RU18a, RU18b]. kinematical [HLMV21]. kinematical/dynamical [HLMV21]. kinesin [Bol21]. know [FV11].

Knowledge [Lin22, All23, Bei12, Dun22, GS18, KP15, KB11, Mil14]. knowledge-what [GS18]. Kochen [Acu21]. Königsberg [Räz18]. Kuhn [GS22a, GS22b].

lab [KM16b]. labor [Bed19]. labour [Lew19]. laden [KK21]. ladenness [Pol21]. Lakatos [Mar20, Dim20a, Mar19]. Lakatosian [De 23]. Lanford [Val14]. language [Bir12, Lut12]. late [Gio18, PES22]. law [Fis24]. Laws [Fri16, SGH21, Bar19, CPW20, HLMV21, Hir21, ILP21, Kri20, Kri21, Lin20, Moh22]. layering [Fil21]. learn [KM16a]. learning [BS21, Kar21, Mül24, Mus22, SR22, Yee23]. legacy [Sto21a, Sto21b]. legitimacy [Hil22]. legitimate [Int15]. length [Gio23]. lenses [San21]. lessons [Kat23, Pin15]. letter [IR17, IR18a, IR18b]. level [Sch22a, Woo22]. levels [Hoc22]. Levin [Ala12]. LHC [Mät22, Pas19]. liberalization [Sar13]. libertarian [Hei14]. life [Jac22, Mei19]. light [PES22]. limit [BB15]. limitations [Bet15]. limits [Gli21, Gut20, Mul18, Pow12]. lineage [Dup11]. linear [NNW19]. Linnaeus [PS18]. literally [SW22]. living [RU18a, RU18b]. Local [Vir21, EE14]. locality [San14]. lockdowns [Dou23]. logic [BP21a, BDPP15, BC22]. logical [BC22]. look [AD14, Fri15, GS16, Maz14, Pol24]. Looking [Hil22]. loop [Le 20]. Lorentzian [Lin21]. Lost [HM21]. lottery [NP18, Nor18a, Nor18b]. Ludwig [RE24].

Machine [SR22, Yee23, Kar21, Mus22, NP18, Nor18a, Nor18b]. machines [SOH21]. macroeconomics [Hen18]. magic [Vic12]. magnets [KL16, KL23]. make [Maz14, PW18]. Making [Pes11, SD21, Str21, Hea15, Lau18, Mon17, Yee21]. malfunction [BvE21]. Malfunctions [Cas17]. manufacture [BL15]. map [Men20]. marginal [LC21]. Markovian [GS13]. Martí [Häg22]. Masking [Suá23]. material [FN18, Nor11, RE24]. materiality [Lud23]. mathematical [Bak21, Fil21, Han16, Pan24, Pov20]. mathematics [Cel18, De 23, Mac21b, PES22, Räz18, Rov17, VG22]. matter [EG22b, Fee22, OC11, VDE17]. matters [EWVTL24, WG23]. maybe [Lóp23]. McMullin [Sch17]. McMullin-style [Sch17]. mean [Kro21]. meaning [Gio21, Gio22, Pot19]. means [McL11]. measure [Ang13, Bou19, Fle21a]. Measurement [AF22, KH21, San14, Bas21, Lau23, Luc22, Pet18, San12, Ves21, WB22]. measurements [Boc22, Mus22]. Measuring [Mor21, Run21]. mechanical [KG14, SOH21]. mechanics [Bar11, Dar22, Esf13, ET20, GL24, Lau14, Lau23, LL18, McC20, OC11, Toa21, VDE17, Wal23]. mechanism [Au21, Bar22, IW12]. Mechanisms [CK23, CPW20, Bec12, IW12, PG20, Kri20, Kri21]. Mechanistic [PG20, Wri12, vE15, Bae12, Bae22, Dra17, HS19b, KG14, LD21, Zac22]. mediators [MM18]. medicine [BCVV21, BC19, Dra17, GS13, Sha21, Sol11b]. medicines [Lal23]. meet [Acu21]. meets [RP14, Sha18]. memory [And23]. mental [Dim20b, Sal17]. Mercury [Nor11]. Mermin [HS21]. meta [BCVV21, Str21, THP23]. meta-analysis [BCVV21]. meta-hypotheses [Str21]. meta-method [THP23]. metaphorical [SW22]. Metaphors [VN21]. Metaphysical [GL24, All23, Mas21b, Mor19]. metaphysically [Lin20]. metaphysics [Fin21, GL21, Gio18, Jak20, Jak21a, Lóp23, Reu17]. method [Sch15, SM23, THP23]. Methodological [Pet18, All23, Hei14]. Methodology [Fin21, Han19, Jus12, LT21, Odd24]. methods [AJ17, GPM23, Gho23]. Meyerson [Gio18]. Michelangelo [Rov17]. Micro [Sch22a]. Micro-level [Sch22a]. microbiome [DS23]. microstructural [Häg22, HM19]. might [KM16a, SH19]. Mill [Sch15]. mind [Pen21, Sch13]. Minimal [BM23]. minimalist [VDE17]. miracle [Ala12]. Miracles [LR22, Spr16, Sus21]. Misconceptions [Rey21]. Misconstrued [Han22]. misinformation [Yee23]. misrepresentation [PS16]. missing [Bed19]. mistaken [Mac21a]. mitigation [Con21]. mitochondrial [Lew19]. modal [Bar19, HLMV21, Knu21]. modalities [CS23]. Modality [Pov20]. Model [Cha20, JB21a, KL16, Ant21, BM23, Fat23, Fri15, Han23, Mon13, New14, O’L23, RM16, Sch22a, Spr13, Sta12, VFHS23, KL23]. Model-groups [Cha20]. Modeling [FEG19, MP17, SR13, Woo22, CS23, Det24, Dou23, Int15, Jac22, Knu21, Lis17, NNW19, Pes11, SD21, HM21]. modellers [Fum16]. Modelling [Win22, Anz21, JB21a, Kar21, Lam21, PUB22, SS15]. models [Bet15, BP21b, Boo20, CK23, EG22a, Fis24, FN18, Fum16, GS13, Han16, Hea15, HS21, Kar21, KP15, KB11, Kos17, Leo19, MM18, PW18, Pet18, Pex14, Pol21, VFHS23]. modern [PS18]. modes [MN16]. modified [RU18a, RU18b]. Modular [SB15]. molecular [BDPP15, BC22, BBŠ21, Sei20, Zac22]. molecules [CH23]. MOND [Kas23]. monist [Moh22]. monogenomic [Dup11]. Moral [PUB22]. Moralisation [Lal23]. Mori [tV22]. motility [Bol21]. Motivationalism [Luc23]. movers [Lar22]. much [Pet16, Sch15]. multi [GS13, Gua23, HR18, Woo22]. multi-causal [GS13]. multi-field [HR18]. multi-functional [Gua23]. multi-level [Woo22]. multidisciplinarity [Men20]. Multiple [Aiz13, Kos19, SFG^{+}18b, Slo12, SFG^{+}18a]. muscle [Suá23]. music [CK16]. Musical [CK16]. musing [Dar22]. Must [Acu21].

narrative [Mül24]. Natural [Raa21, Art21, Bar19, Fag22, Gra24, Hir21, Mor21, Pow12]. naturalism [Ala12]. naturalistic [LT21]. naturalized [Jak20, Jak21a]. naturalness [Lud23]. Nature [SGH21, AB15, HLMV21, Moh22, PS18, Suá23]. necessary [Lin20]. Neo [Lin21, DS23, Jak20, KG14]. neo-Hempelian [DS23]. Neo-Lorentzian [Lin21]. neo-mechanistic [KG14]. neo-positivist [Jak20]. Nested [CS23]. network [DS23]. networks [DS23, Mar13, Pan24]. Neumann [Acu21]. neural [Hac22, Pan24]. neuro [Fum16]. neuro-psychological [Fum16]. neurobiological [Joh12]. neuroscience [Hen20]. Newman [Kha15]. Newtonian [Bel19, Fle12]. Nexus [Suá22].

Neyman [Rub20]. niche [Suá23]. Nicolai [Pin15]. nineteenth [Luc22, Hij22]. nineteenth-century [Luc22, Hij22]. No [Fri12, Häg22, ILP21, Ala12, Her14, Lau14, San12, Spr16]. no-conspiracy [San12]. no-go [Lau14]. Non [EE14, LE22, RE22, Cla13, For24, NNW19, Reu17, San14]. non-causal [Reu17]. non-eliminativist [Cla13]. Non-Epistemic [RE22, LE22]. non-linear [NNW19]. Non-local [EE14]. non-locality [San14]. non-platonic [Cla13]. non-spatiotemporal [For24]. nonclassical [WG17]. noncontextual [HS21]. nonlocality [Lau18]. Nor [Lop24]. normal [LWW21]. normativity [Sto21a, Sto21b]. norms [Höl21, Mül24, Ruy20]. Norton [Fle12]. Note [Teh15, Her14]. Nothing [Gio21, CV20, Gio22, Sha18, Sha22]. notion [Bar11, EW22]. Novel [SD20, Lar22]. nudges [Hei14]. Numerical [AJ21, AJ17].

Object [Sha18]. objection [Hit12]. objections [Par19]. Objectivity [HH23, vDS21, Eva20, IdMM14, Spr18]. observation [Hel17]. observational [Wer11]. observe [Pas19]. offs [Fre23]. oil [ET20]. Old [Jak20, Fri15, Sag23, Vic12]. one [Bou19, KM16a, McL11]. one-to-one [Bou19]. Ontic [Esf13, Mor19, Wri12]. onto [PES22]. ontological [OÖ21]. ontologies [Fum16, KG14]. ontologists [Bel12]. ontology [Aus20, Egg21, EG22b, For24, Koo21, Lam15, Lop24, VDE17]. operational [El 21]. operationalization [Con18]. operations [NH23]. operator [tV22]. opinions [Hil22]. opposition [Bas21]. options [Luc23]. Organisms [Aus20]. organization [MP17]. oriented [BOS22, VH21]. other [Bei18, Wal23]. ought [Pol24]. our [Str21]. Overcoming [GY22]. Overextension [Sch13]. overlapping [Lop24]. overlooked [Luc23]. Overstating [GPM23].

p [Aut16]. p-values [Aut16]. paleoclimatology [WB22]. paleodiversity [Boc22]. pandemic [KH21, LC21]. Pandemics [Dou23]. papers [RTT19]. paradigm [Sol11b]. paradigms [BV19]. paradox [CV20, Lae11, Sha18, Sha22, VSF17]. parameter [San14]. paribus [Spo14]. partial [Luk21, Roc16]. particles [Big20]. Partitions [FS21]. past [BS20, Lóp19, Vez17]. patchwork [Tay23a]. Patchworks [NH23]. paternalism [Hei14]. path [Ben17]. paths [Slo12]. Pattern [Hel17]. Pearson [Rub20]. Pedro [Wro14]. Peirce [BP21a]. Pemberton [Kri20, Kri21]. Peres [HS21]. perihelion [Nor11]. permeate [BP15]. Perrin [CH23]. persist [LR22]. Perspectival [Eva20, Ruy20, VB21, San21]. perspective [Bou24, FM18, HM20, KH21, Luc22, MM18, PES22]. perspectives [Boo20, Hil22, MM16]. Perspectivism [Bsc20, Pot20, Ber20, Kha22, Mas21b, Bak20, Fel21]. pessimistic [Bog21]. pharmacology [LOP18]. phenomena [Col20]. phenomenological [Ber20, Kha22]. phenomenological-hermeneutical [Kha22]. phenomenon [Bol21, Pot19]. phenomenotechnique [Sim18]. philosophers [BOS22]. philosophical [Bec12, BC19, BV19, Dar22, VSF17]. Philosophy [Mäk16, Sto21b, Aam22, ACBB11, Auf11, Ber19, Cab21, CC23, Dim20a, EW22, GAD^{+}21, GY14, GY22, HS12, Jak21b, Jus12, Kle21, Kor18, Lau14, Lud23, Lut12, Mac21c, MM16, Pin15, Rey21, SS15, SOH21, Spr13, WG23, You22, dRK21, Sto21a]. Physic [Dol18]. physical [De 19, Dim20b, Eva16, Lóp23, Luc23]. physicalism [Bae22, DiF18, Ero12]. Physicist [HTS24]. physicists [JFEH19]. physics [Cro15, DG23, Dim20b, Fri12, Gio21, Gio22, Hub21, JFEH19, Kin20, Kob23, Lan19, Lin20, Pex16, SR22, Gio18]. place [Fri12]. Placebo [JB21b]. planetary [LR24]. platonic [Cla13]. platonism [Rov17]. player [Fre23]. Playing [Mar13]. Plenitudinous [Sha22]. Pliability [Tam14]. pluralism [CK16, KM16a, Rey21, VB21, SW21]. pluralist [De 23]. Pluralistic [Ero12]. plurality [EWVTL24]. Poincaré [Dan21]. point [Gio21, Gio22]. point-coincidence [Gio21, Gio22]. points [LR24]. polarization [OW18]. policies [LT21]. policy [Bir21, Bir23, Mon17, Pae24, You22, KP15]. policy-relevant [Pae24]. political [FRT24]. politics [EG22b]. polysemy [Şer17]. Pope [Dan21]. population [Rub20]. populations [De 23]. position [FD21]. positivist [Jak20]. possibilistic [Bet15]. possibilities [Kat23]. possibility [All22, Bri14, IK21]. possibility- [Bri14]. possible [Ang13]. possibly [Kos17]. postgenomic [Bel22]. Potentiality [DG23]. potentially [FX21]. power [BCVV21, Räz18, Sal17]. Powers [All22, Koo21, GL21, ILP21]. practical [PUB22, SS15]. practice [ACBB11, Bas21, BOS22, GAD^{+}21, Kle21, SH19, VG22, Fre20]. practice-based [GAD^{+}21]. practice-oriented [BOS22]. practices [Ans21, Sol11a]. pragmatic [Can22, El 21, Fri22, GJB23, KK21, McL11, Mil14, Pet16, SM23]. praise [Dou23]. precision [BC19]. predicting [SR22]. prediction [Bet15]. predictions [Mus22]. predictive [Mus22]. Predictivism [Fat23, Fri15]. preferences [KM16b]. pregnancy [Fin21]. Presentism [RP14, Sen17]. presumed [AJ17]. prevalence [NNW19]. Price [SGH21]. primary [Bel19]. Primitive [Lam15, Bel12]. principal [Fis21, LWW21, WW20]. principle [Aut16, BP21a, Fis21, Hen18, Lan19, OC11, Wil18, LWW21, WW20]. priors [Mas21a]. probabilistic [Roc12, Sob20, Spr16]. probability [Ber19, Par19, Roc16, Roc17, She20, WG17, tV22]. Probing [El 21]. problem [All22, Bel19, Boc22, CK23, Hac22, Kas23, Lau23, MN16, Sag23]. problem- [MN16]. problems [IdMM14, Jak20]. procedural [BC22]. Procedure [De 21]. Process [DL22, Aus20, KH21, Mei19]. processes [Bir12]. processual [BC22]. produce [Bei12]. profit [IdMM14]. profit-driven [IdMM14]. programmes [Cha20]. progress [JB21a]. Progressive [Dun22]. project [SR22]. projection [tV22]. projections [Hea15]. proliferation [Sim18]. propensities [DG23]. propensity [Suá14]. Proper [Val16]. property [RU18a, RU18b]. property-cluster [RU18a, RU18b]. Prospects [Bet15]. protests [Mar19, Mar20]. proving [Pan24]. Proxy [WB22]. psychiatric [Bas21, Fel21, GJB23]. psychological [Fee20, Fee22, Fle21b, Fum16, Joh12, Win22]. psychology [Hen20, She21]. psychometric [Spe18]. Public [IN22]. publishing [Dun22]. pure [Bar11, CC23, PES22]. purification [Sim18]. Purifying [PES22]. purpose [BP21b, LE22]. pursuing [MS23]. pursuit [Han23, McL11]. Pursuitworthiness [Vir23]. Putting [Men20].

QBism [Gli21]. qualia [Far20]. qualitative [Pol24, Zah23]. qualities [Bel19]. quality [Fee22, Dra17]. quanta [Nor11]. quantification [Luc22]. quantity [Au21]. Quantum [Bel12, Egg21, Lop24, Shr14, Esf13, ET20, EG22b, Fie22, GL24, Jak23b, Koo21, KG14, Lam15, Lau14, Lau23, Le 20, Lóp19, LL18, McC20, Odd24, OC11, OÖ21, Old24, Pin15, Toa21, Vic12, Wal23]. quantum-realist [EG22b]. question [Wan24]. question-centered [Wan24].

R [Pen21]. racist [Spe18]. Radical [dO22]. ramseyfications [Ueb11]. rational [Dim20a]. rationalistic [Gio18]. razor [LR20]. Reactive [Fag22]. Reactivity [JB21b, Kos22, MZG24, Run21, Zah23, Fee22]. real [Hit12]. Realism [HM20, AB15, Ben17, Ben21, Bog21, DG23, Díe18, Dop14, El 21, Esf13, FV11, Gio18, Gli21, HTS24, Jak23a, Kro11, Mor19, Ney23, Ruy20, Saa22, Tam14]. realist [EG22b, Kha22]. realistic [DG23, Fum16]. Realities [EG22b]. reality [CH23, Gio23, Jak23a]. Realizability [Kos19]. realization [Aiz13]. really [Lae11]. reappraisal [Art21, Sar15]. reasonable [LWW21]. reasoning [GY22, VH21, Vez17, Win22]. reasons [Lóp23, Wal23]. Recharacterizing [Col20]. recommendations [McL11]. reconsidered [Ban17]. Reconstituting [Bol21]. Reconstructing [Ben17, Dim20a]. Reconstructions [Odd24, Dim20a]. redefine [Tun22]. reduction [Cro15]. reductionism [Rae18]. reductive [Bae22]. reference [HM20, Kro11]. refinement [HS19b]. reflections [Mas21b]. reflexive [MK21]. Reflexivity [Nor22]. reframing [Mus22]. regular [Par19]. Regularity [How17]. regulatory [LT21]. reification [DL22]. reimagining [Lud23]. Reining [Sar13]. related [Bei18]. relation [CCS22, Dim20a, KG14]. Relationalism [VDE17]. relationism [Maz16]. relations [Mac21b, Vas20]. relationship [Joh12, Luc23]. relative [CDR23]. relativism [Fel21]. Relativistic [Dob19, Aam22]. Relativity [Kas23, Sus21, Lin21, Val16, Vas20]. relevance [Bae12, San12]. relevant [Pae24]. reliability [Cab21]. renormalization [Sul19]. reorienting [You22]. Repeated [Rub20]. replicability [Gut20]. replication [Fle21a, Fle21b]. replications [Str21]. Reply [Kri20, Kri21, Häg22, Joh15a, Joh15b, Kha15, LR22, Par19, Roc16, VNT22]. Representation [Ves21, BP21b, Pes11, Ruy20]. representational [Anz21]. representations [Hac22]. representing [FN18]. reproducibility [FRT24, Hen20, HW21]. research [Au21, BP21a, BV19, BOS22, Cha20, DL22, FD21, IdMM14, Kat23, Kos22, Pol24, Reu20, SH19, vDS21]. researcher [FD21]. researchers [EWVTL24]. resistance [McL11, Tam14]. resources [BS21]. Respecting [WR23, Fel21]. response [Bir21, Bir23, Par21, SM23]. responses [Rub20]. responsibility [Pol24]. responsible [dMMI23]. Resuming [Näg22]. Rethinking [She21, DL22]. reversal [Lop21]. review [SR22]. revisionary [Fri22]. Revisiting [Zac22]. revolution [Koo21]. Richard [Her14]. rise [TV17]. risk [Con21, EW22, Hen21, HH23, Kar21, Pae24]. Roads [Lóp19]. Robust [Bir12, Gur20]. Robustness [FS21, Lis17, Han16]. role [And23, Boo20, Fle21b, HS19a, Hea15, LC21, Mon13, Odd24, Spr13, Str21, Suá23]. roles [BOS22]. room [HW21]. Rovelli [Teh15]. rules [Fle19].

s [Lau23]. same [Mol22, Ros13, Rub20]. sampling [Rub20]. San [Wro14]. sans [GLB15]. scanner [KM16b]. scheme [Vir23]. scholarly [Dun22]. Science [Bir21, Mil14, RR24, And23, ACBB11, Ans21, Auf11, Aus17, Bak21, BvE21, Bar22, BP21a, BL15, BV19, BOS22, BS21, Cab21, CC23, CK16, Dan21, Dim20a, EW22, Fle21b, GPM23, GL21, GAD^{+}21, GY14, GY22, Han20, HS12, Hil22, Hoc22, HH23, IN22, Jak21a, Jak21b, Jak23b, Jus12, Kat23, Kid21, Kor18, LR24, LC21, LT21, Lus22, Lut12, Mac21c, MM16, MP17, Mas21b, McL11, MS23, Mon13, Mon17, Nor11, Pae24, Pet21, Pol21, Rey21, Sal17, Sha21, SOH21, Str21, VN21, VNT22, Vir21, WG23, Yee23, You22, dMMI23, dRK21, vE15, Bir23]. sciences [BG21, De 19, Dup11, Fil21, IW12, LC21, MZG24, Pet18, SW21, VH21]. Scientific [Ban17, Ber20, DM23, Koz23, OW18, Bec12, Boo20, BOS22, BFŠS19, Can22, Cha20, Chu22, Col20, Dop14, FD21, Gli21, Han19, HTS24, Jak21b, JB21b, KM16a, KK21, Lud23, LE22, MZG17, Nap21, New14, Pol21, Pol24, Ruy20, SR22, SD21, Tun22, THP23, VH21, Fel21, Fre20]. scientism [THP23]. scientists [GM22, GY14, GY22]. search [Fri13, Jac22, JFEH19, Kos17]. searches [Kob23]. Second [Cab21]. securing [Sta12]. see [EWVTL24]. Selected [RTT19]. selection [Art21, Fat23, Pow12, RM16, Spr13]. selective [Bog21]. Self [MK21, GL21, Hea21]. self-identity [Hea21]. Self-reflexive [MK21]. Semantic [Fil21, Le 12]. sense [Bog19, Bog22, Lau18, Pes11, Yee21]. senses [SD21]. sensitive [Höl21]. seriously [GS22a, GS22b, Han23, Kos17]. set [Trl22]. Setting [MP23]. Shackel [CV20]. sheds [PES22]. shifting [BV19]. shock [Pex16]. Should [GJB23, GS13, Lóp23, GM22, Iva19, Mon17, SH19, Tun22]. show [Pet16]. side [Due23]. sides [Mol22]. signals [Suá23]. Significance [Aut16, Gom21, VSF17]. silence [Dol18]. simple [Lar22]. Simplicity [RM16]. Simplified [MM18]. simulations [Bei12, Bei18, Bog19, Bog22]. single [RW11]. single-case [RW11]. Sisyphean [MS23]. situated [VG22]. skepticism [BL15, Fri12]. sky [Wal23]. slid [Ros13]. slippery [Ros13]. slope [Ros13]. Social [Gua23, Han20, BG21, Ben21, Fee20, GJB23, Hil22, Jak23b, JB21b, KM16b, LC21, MP17, Mül24, PG20, Pet18, Pol24, SW21, SOH21]. Socially [dMMI23]. societies [Anz21]. socio [KL16, KL23]. socio-economic [KL16, KL23]. software [HW21]. solution [Lae11, Pov21]. solutions [AJ17]. solving [MN16]. some [Kat23, Luc23]. sophisticated [Tam14]. south [Shr14]. space [Maz14]. spaces [MZG17, WG17]. spacetime [DdR18, For24, Slo12]. spacetimes [Dob19]. spatial [Rya24]. spatiotemporal [For24]. special [BC17]. specificity [Bou19]. Specker [Acu21]. speculation [Egg21]. spillovers [Lin22]. sponsorship [Reu20]. spread [JB21a]. Sprenger [Her14]. square [HS21]. stability [Han16, Lam21]. Stabilization [Pot19]. stance [LD21]. standards [LWW21]. state [BS20]. states [Bel12]. static [TB21]. statistical [Kuo12, McC20]. status [FRT24, Fie22, HLMV21]. Stein [VSF17]. stem [Suá23, Cab21, Lus22]. Step [De 21]. Stephan [Her14]. Still [Maz16, Häg22]. stone [Rov17]. Stop [Lau18]. Stopping [Fle19]. Strategies [Sta12, KG14]. strategy [Mac21c]. String [Le 20, HM21]. strong [Hen21, Sei20]. Structural [AB15, Ben17, Ben21, Esf13, FV11, Lam21, Mor19, NNW19]. structuralism [Eva16, Slo12]. Structure [Höl21, AB15, PS18, Sei20, Win22, WR23]. Structure-sensitive [Höl21]. structures [BBŠ21, SR13]. Structuring [Kob23]. struggle [Gio21, Gio22]. student [Lus22, Mar19, Mar20]. students [BS21, Cab21, GAD^{+}21, VH21, You22]. studies [Mät22]. study [Bak21, Lus22, Pol24]. style [FEG19, Sch17]. subject [Fee22]. subjective [Fis21, WW20, Spr18]. substance [Aus20]. substantivalism [Maz16]. substitute [Han16]. Success [Hei14, Fil21, GY16, SD21, Wan24]. suggest [Gho23]. superiority [AJ17]. supervenience [Woo22]. support [Lae11, Roc12]. surprise [Koz23]. Surrealistic [Sol17]. Sus [LR22]. sustainability [Sto21a, Sto21b]. symmetries [Gom21, Lóp23, Luc23]. Symmetry [Luk21, Par19, Fri13, Lop21]. symptom [Bas21]. symptom-based [Bas21]. Synchronic [Big20, Sar15]. synthesis [PS18, GS22a, GS22b]. Synthetic [Kos17, IK21]. system [Fle12, LLW13, LR24]. systemic [JFEH19]. systems [AJ21, Dow19, Fri22, Han16, KL16, KL23, Kos17, MN16, Mac21b, Woo22].

’t [BB15]. Tacit [Sol11a]. Tacking [Sch22b]. Tahko [HLMV21]. takes [Le 12]. Taking [Han23, Kos17, GS22a, GS22b]. talk [Iva19]. taxonomy [Con18]. teach [BS21]. Teaching [GY14, Jak21b, WG23, BOS22, GAD^{+}21, GY22, Rey21, VH21, dRK21]. teleological [Bir12]. teleology [Cas17]. tell [ET20, Şer17]. templates [KL16, KL23, Lin22]. Temporal [Aam22, Dow19, Far20]. terms [Raa21]. Testability [DM23]. testimonial [Cab21, Höl21]. testing [Aut16, Bar18, PUB22]. tests [Bae12, Fee20, Spe18]. textbooks [Lop24]. their [SFG^{+}18a, SFG^{+}18b]. themselves [EWVTL24]. Theorem [Pan24, GH23, Lau18, Val14]. Theorems [Bac13]. theoretic [Bou24, Lau23]. Theoretical [Sch17, WR23]. theories [Acu21, Bai13, BB15, DdR18, For24, Fri13, Lóp19, Şer17, Slo12, Suá14]. Theory [BFŠS19, Kro11, RE22, VNT22, AD14, Bar18, De 19, Díe18, Dop14, Dra17, Eva16, Han16, Han22, HM21, Kid21, Lam15, Le 20, MZG17, Nor11, Odd24, OÖ21, Reu17, Sob20, Trl22, Val16, Vic12, VG22, Wel11].

Theory-choice [BFŠS19]. Theory-dependence [Kro11]. thermodynamic [BS20]. Thesis [De 21]. Thick [AF22, SD20]. thin [ILP21]. things [RU18a, RU18b]. Thinking [IW12]. third [Nap21]. Thought [Hit12, Cla13]. Three [HS21, Lop21, Ney23, Fum16, Mät22, De 21]. Three-Step [De 21]. time [Aam22, Adl22, Dol18, Fie22, Lóp19, Lop21, Maz14, Val16, Wer11]. time-reversal [Lop21]. tipping [LR24]. Token [DiF18]. too [Pet16]. tool [Kos22]. toolbox [BDPP15]. toolkits [Rey21]. tools [Sol11a]. topical [Mas21b]. topology [Rya24]. tossed [How17]. total [Aut16]. tractability [Lis17]. trade [Fre23]. trade-offs [Fre23]. tradition [Ber20]. traditional [Sha21]. trait} [Kõi23]. trajectories [Sol17]. transdisciplinarity [KM16a]. transfer [Can22, Lin22, Sto21a, Sto21b]. transient [BFŠS19]. transitivity [Roc12, Roc17]. treatment [SM23]. triad [Pov21]. Triangulation [KM16b]. Trilemma [Díe18]. trouble [Hen20]. tuning [Bar18, Fri15]. tunnelling [Fie22]. Turing [De 21]. turn [Fri22, FN18, Kid21]. Two [Har19, Mol22, Bae12, Bed19, Fre23, Jak23b, RR24, Lóp23]. two-variable [Bae12]. two-way [RR24].

UK [Bir23, Bir21, Hea15]. uncertainties [GJB23]. uncertainty [SH19]. unconceived [Kas23]. underdetermination [AD14, Rya24, Woo22]. underdetermined [Wal23]. undergraduate [Kle21, Lus22]. Understanding [Bec12, Hub21, SW22, VJ19, tV22, Ban17, New14, Pot11, Tay23b, Van14]. understandings [EWVTL24]. Unexpected [Old24]. Unexplained [JP21]. unification [Fry16]. unificationism [Ban17]. Universality [Sul19]. universalization [Bel19]. universe [Aam22, Gur20, Lin21]. Unmoved [Lar22]. unsound [Jak23b]. upward [Ben17]. Using [MZG17, Mon17, BS21].

validity [Fee20]. valuable [FX21, Lus22]. value [Bet13, Fis24, HH23, KK21, MS23, Pol21, Joh15a, Joh15b]. value-free [HH23]. value-laden [KK21]. value-ladenness [Pol21]. Values [Kar21, PW18, PUB22, RE22, Au21, Aut16, Hil22, Int15, KP15, LR24, LE22, Mil14, RR24]. valve [FN18]. variable [Bae12, HS21]. variables [Acu21, BD13]. Variation [Bou19]. Varieties [Hir21, PS16, Fag22]. Variety [Lan20, Vez17, Au21]. Variety-of-evidence [Vez17]. version [RU18a, RU18b]. versus [RW11, SM23, Sha22, VG22]. very [Lar22]. viability [DM23]. view [CC23, Fle12, HLMV21, Lau23, LFL16, LE22, Mei19, Le 12]. views [HTS24]. vindicate [Joh15a, Joh15b]. virtual [Lan19]. virtue [Kid21, Tun22]. Visibility [Ans21]. vs [Boc22, Luc23, Sar15, She20].

W. [Pen21]. Walrasian [Han16]. wanted [FV11]. warranted [Kro11]. Water [HM19, Häg22]. wave [Bar11, HR18, Ney23]. wave-function [HR18]. way [RR24]. weaker [Roc12]. web [SOH21]. Weintraub [Par21]. Weldon [Pen21]. welfare [VB21]. were [FV11]. Who [Pol24, Kit19]. whys [dRK21].

Wieten [Kri20, Kri21]. Williamson [Par21]. within [EWVTL24, EG22b, KL16, KL23, Lus22, McL11, Spr13, Web22]. without [DdR18, Egg21, GY16, Lop24, Sus21, Wri12]. Wittgensteinian [PES22]. Woodward [Näg22]. work [Lan19]. worker [vDS21]. working [Ans21]. worlds [Bar11, Bet15]. worry [Mon17]. worse [JB21b]. worth [MS23]. worthy [SD20]. wrong [Reu20].

Zalabardo [Roc16]. Zhuāngzı [Chu22]. Zwanzig [tV22].

References

Aames:2022:TBR

[Aam22] Jimmy Aames. Temporal becoming in a relativistic universe: causal diamonds and Gödel’s philosophy of time. European Journal for Philosophy of Science, 12(3):??, September 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-022-00471-z.

Arenhart:2015:SRN

[AB15] Jonas R. Becker Arenhart and Otávio Bueno. Structural realism and the nature of structure. European Journal for Philosophy of Science, 5(1):111–139, January 2015. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-014-0100-y.

Ankeny:2011:IPS

[ACBB11] Rachel Ankeny, Hasok Chang, Marcel Boumans, and Mieke Boon. Introduction: philosophy of science in practice. European Journal for Philosophy of Science, 1(3):??, October 2011. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-011-0036-4.

Acuna:2021:MHV

[Acu21] Pablo Acuña. Must hidden variables theories be contextual? Kochen & Specker meet von Neumann and Gleason. European Journal for Philosophy of Science, 11(2):??, June 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-021-00347-8.

Acuna:2014:ALE

[AD14] Pablo Acuña and Dennis Dieks. Another look at empirical equivalence and underdetermination of theory choice. European Journal for Philosophy of Science, 4(2):153–180, May 2014. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-013-0080-3.

Adlam:2022:DBT

[Adl22] Emily Adlam. Determinism beyond time evolution. European Journal for Philosophy of Science, 12(4):??, December 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-022-00497-3.

Alexandrova:2022:DMW

[AF22] Anna Alexandrova and Mark Fabian. Democratising measurement: or why thick concepts call for coproduction. European Journal for Philosophy of Science, 12(1):??, March 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-021-00437-7.

Aizawa:2013:MRC

[Aiz13] Kenneth Aizawa. Multiple realization by compensatory differences. European Journal for Philosophy of Science, 3(1):69–86, January 2013. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-012-0058-6.

Ardourel:2017:PSA

[AJ17] Vincent Ardourel and Julie Jebeile. On the presumed superiority of analytical solutions over numerical methods. European Journal for Philosophy of Science, 7(2):201–220, May 2017. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-016-0152-2.

Ardourel:2021:NID

[AJ21] Vincent Ardourel and Julie Jebeile. Numerical instability and dynamical systems. European Journal for Philosophy of Science, 11 (2):??, June 2021. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-021-00372-7.

Alai:2012:LGN

[Ala12] Mario Alai. Levin and Ghins on the “no miracle” argument and naturalism. European Journal for Philosophy of Science, 2(1):85– 110, January 2012. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-011-0028-4.

Allen:2022:PHP

[All22] Sophie R. Allen. Powers and the hard problem of consciousness: conceivability, possibility and powers. European Journal for Philosophy of Science, 12(2):??, June 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-022-00451-3.

Allzen:2023:AMC

[All23] Simon Allzén. Against methodological continuity and metaphysical knowledge. European Journal for Philosophy of Science, 13 (1):??, March 2023. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-022-00505-6.

Andonovski:2023:AGS

[And23] Nikola Andonovski. Autonoesis and the Galilean science of memory: Explanation, idealization, and the role of crucial data. European Journal for Philosophy of Science, 13(3):??, September 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-023-00548-3.

Angner:2013:IPM

[Ang13] Erik Angner. Is it possible to measure happiness? European Journal for Philosophy of Science, 3(2):221–240, May 2013. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-013-0065-2.

Anscomb:2021:VCC

[Ans21] Claire Anscomb. Visibility, creativity, and collective working practices in art and science. European Journal for Philosophy of Science, 11(1):??, March 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-020-00310-z.

Antoniou:2021:WDM

[Ant21] Antonis Antoniou. What is a data model? European Journal for Philosophy of Science, 11(4):??, December 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-021-00412-2.

Anzola:2021:CRE

[Anz21] David Anzola. Capturing the representational and the experimental in the modelling of artificial societies. European Journal for Philosophy of Science, 11(3):??, September 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-021-00382-5.

Artiga:2021:BFN

[Art21] Marc Artiga. Biological functions and natural selection: a reappraisal. European Journal for Philosophy of Science, 11(2):??, June 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-021-00357-6.

Au:2021:EVQ

[Au21] Yin Chung Au. Epistemic values of quantity and variety of evidence in biological mechanism research. European Journal for Philosophy of Science, 11(2):??, June 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-021-00369-2.

Aufrecht:2011:CDC

[Auf11] Monica Aufrecht. The context distinction: controversies over feminist philosophy of science. European Journal for Philosophy of Science, 1(3):??, October 2011. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-011-0031-9.

Austin:2017:EDS

[Aus17] Christopher J. Austin. Evo-devo: a science of dispositions. European Journal for Philosophy of Science, 7(2):373–389, May 2017. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-016-0166-9.

Austin:2020:OAB

[Aus20] Christopher J. Austin. Organisms, activity, and being: on the substance of process ontology. European Journal for Philosophy of Science, 10(2):??, May 2020. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-020-0278-0.

Autzen:2016:STP

[Aut16] Bengt Autzen. Significance testing, p-values and the principle of total evidence. European Journal for Philosophy of Science, 6(2): 281–295, May 2016. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-016-0138-0.

Bacciagaluppi:2013:ITE

[Bac13] Guido Bacciagaluppi. Insolubility theorems and EPR argument. European Journal for Philosophy of Science, 3(1):87–100, January 2013. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-012-0057-7.

Baetu:2012:FMD

[Bae12] Tudor M. Baetu. Filling in the mechanistic details: two-variable experiments as tests for constitutive relevance. European Journal for Philosophy of Science, 2(3):337–353, October 2012. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-011-0045-3.

Baetu:2022:MGR

[Bae22] Tudor M. Baetu. A mechanistic guide to reductive physicalism. European Journal for Philosophy of Science, 12(4):??, December 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-022-00489-3.

Bain:2013:EEF

[Bai13] Jonathan Bain. Emergence in effective field theories. European Journal for Philosophy of Science, 3(3):257–273, October 2013. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-013-0067-0.

Baker:2020:GIP

[Bak20] Kane Baker. Giere’s instrumental Perspectivism. European Journal for Philosophy of Science, 10(3):??, October 2020. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-020-00296-8.

Baker:2021:BGC

[Bak21] Alan Baker. Bipedal Gait Costs: a new case study of mathematical explanation in science. European Journal for Philosophy of Science, 11(3):??, September 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-021-00411-3.

Bangu:2017:SEU

[Ban17] Sorin Bangu. Scientific explanation and understanding: unificationism reconsidered. European Journal for Philosophy of Science, 7(1):103–126, January 2017. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-016-0148-y.

Barrett:2011:EPW

[Bar11] Jeffrey A. Barrett. Everett’s pure wave mechanics and the notion of worlds. European Journal for Philosophy of Science, 1(2):277– 302, May 2011. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-011-0023-9.

Barnes:2018:FTC

[Bar18] Luke A. Barnes. Fine-tuning in the context of Bayesian theory testing. European Journal for Philosophy of Science, 8(2):253– 269, May 2018. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-017-0184-2.

Bartels:2019:EMF

[Bar19] Andreas Bartels. Explaining the modal force of natural laws. European Journal for Philosophy of Science, 9(1):??, January 2019. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-018-0225-5.

Barman:2022:FME

[Bar22] Kristian González Barman. Fictional mechanism explanations: clarifying explanatory holes in engineering science. European Journal for Philosophy of Science, 12(2):??, June 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-022-00468-8.

Basso:2021:MCP

[Bas21] Alessandra Basso. From measurement to classificatory practice: improving psychiatric classification independently of the opposition between symptom-based and causal approaches. European Journal for Philosophy of Science, 11(4):??, December 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-021-00424-y.

Bouatta:2015:EGT

[BB15] Nazim Bouatta and Jeremy Butterfield. On emergence in gauge theories at the ’t Hooft limit. European Journal for Philosophy of Science, 5(1):55–87, January 2015. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-014-0098-1.

Brzovic:2021:EHC

[BBŠ21] Zdenka Brzović, Vito Balorda, and Predrag Šustar. Explanatory hierarchy of causal structures in molecular biology. European Journal for Philosophy of Science, 11(2):??, June 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-021-00380-7.

Bhakthavatsalam:2017:WSA

[BC17] Sindhuja Bhakthavatsalam and Nancy Cartwright. What’s so special about empirical adequacy? European Journal for Philosophy of Science, 7(3):445–465, October 2017. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-017-0171-7.

Boniolo:2019:CIP

[BC19] Giovanni Boniolo and Raffaella Campaner. Complexity and integration. A philosophical analysis of how cancer complexity can be faced in the era of precision medicine. European Journal for Philosophy of Science, 9(3):??, October 2019. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-019-0257-5.

Boniolo:2022:LEM

[BC22] Giovanni Boniolo and Raffaella Campaner. The logic of explanation in molecular biology: historical–processual and logical– procedural aspects. European Journal for Philosophy of Science, 12(1):??, March 2022. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-022-00445-1.

Berchialla:2021:PMA

[BCVV21] Paola Berchialla, Daniele Chiffi, Giovanni Valente, and Ari Voutilainen. The power of meta-analysis: a challenge for evidencebased medicine. European Journal for Philosophy of Science, 11 (1):??, March 2021. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-020-00321-w.

Baumgartner:2013:IIV

[BD13] Michael Baumgartner and Isabelle Drouet. Identifying intervention variables. European Journal for Philosophy of Science, 3(2): 183–205, May 2013. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-012-0063-9.

Boniolo:2015:ALT

[BDPP15] Giovanni Boniolo, Marcello D’Agostino, Mario Piazza, and Gabriele Pulcini. Adding logic to the toolbox of molecular biology. European Journal for Philosophy of Science, 5(3):399–417, October 2015. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-015-0118-9.

Bechtel:2012:UEA

[Bec12] William Bechtel. Understanding endogenously active mechanisms: a scientific and philosophical challenge. European Journal for Philosophy of Science, 2(2):233–248, May 2012. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-012-0046-x.

Bedessem:2019:DCL

[Bed19] Baptiste Bedessem. The division of cognitive labor: two missing dimensions of the debate. European Journal for Philosophy of Science, 9(1):??, January 2019. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-018-0230-8.

Beisbart:2012:HCC

[Bei12] Claus Beisbart. How can computer simulations produce new knowledge? European Journal for Philosophy of Science, 2(3):395– 434, October 2012. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-012-0049-7.

Beisbart:2018:CSE

[Bei18] Claus Beisbart. Are computer simulations experiments? And if not, how are they related to each other? European Journal for Philosophy of Science, 8(2):171–204, May 2018. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-017-0181-5.

Belot:2012:QSP

[Bel12] Gordon Belot. Quantum states for primitive ontologists. European Journal for Philosophy of Science, 2(1):67–83, January 2012. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-011-0024-8.

Belkind:2019:DNI

[Bel19] Ori Belkind. In defense of Newtonian induction: Hume’s problem of induction and the universalization of primary qualities. European Journal for Philosophy of Science, 9(1):??, January 2019. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-018-0241-5.

Bellazzi:2022:EPG

[Bel22] Francesca Bellazzi. The emergence of the postgenomic gene. European Journal for Philosophy of Science, 12(1):??, March 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-022-00446-0.

Beni:2017:RUP

[Ben17] Majid Davoody Beni. Reconstructing the upward path to structural realism. European Journal for Philosophy of Science, 7 (3):393–409, October 2017. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-016-0167-8.

Beni:2021:ISA

[Ben21] Majid D. Beni. Inflating the social aspects of cognitive structural realism. European Journal for Philosophy of Science, 11(3):??, September 2021. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-021-00401-5.

Berkovitz:2019:FIP

[Ber19] Joseph Berkovitz. On de Finetti’s instrumentalist philosophy of probability. European Journal for Philosophy of Science, 9(2):??, May 2019. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-018-0226-4.

Berghofer:2020:SPP

[Ber20] Philipp Berghofer. Scientific perspectivism in the phenomenological tradition. European Journal for Philosophy of Science, 10(3): ??, October 2020. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-020-00294-w.

Betz:2013:DVF

[Bet13] Gregor Betz. In defence of the value free ideal. European Journal for Philosophy of Science, 3(2):207–220, May 2013. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-012-0062-x.

Betz:2015:CMC

[Bet15] Gregor Betz. Are climate models credible worlds? Prospects and limitations of possibilistic climate prediction. European Journal for Philosophy of Science, 5(2):191–215, May 2015. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-015-0108-y.

Borg:2019:TCT

[BFŠS19] AnneMarie Borg, Daniel Frey, Dunja Šešelja, and Christian Straßer. Theory-choice, transient diversity and the efficiency of scientific inquiry. European Journal for Philosophy of Science, 9 (2):??, May 2019. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-019-0249-5.

Beck:2021:NFS

[BG21] Lukas Beck and James D. Grayot. New functionalism and the social and behavioral sciences. European Journal for Philosophy of Science, 11(4):??, December 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-021-00420-2.

Bigaj:2020:SDI

[Big20] Tomasz Bigaj. Synchronic and diachronic identity for elementary particles. European Journal for Philosophy of Science, 10(3): ??, October 2020. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-020-00298-6.

Birch:2012:RPT

[Bir12] Jonathan Birch. Robust processes and teleological language. European Journal for Philosophy of Science, 2(3):299–312, October 2012. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-011-0043-5.

Birch:2021:SPE

[Bir21] Jonathan Birch. Science and policy in extremis: the UK’s initial response to COVID-19. European Journal for Philosophy of Science, 11(3):??, September 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-021-00407-z. See cor-

rection [Bir23].

Birch:2023:CSP

[Bir23] Jonathan Birch. Correction to: Science and policy in extremis: the UK’s initial response to COVID-19. European Journal for Philosophy of Science, 13(3):??, September 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-023-00546-5. See [Bir21].

Biddle:2015:CSM

[BL15] Justin B. Biddle and Anna Leuschner. Climate skepticism and the manufacture of doubt: can dissent in science be epistemically detrimental? European Journal for Philosophy of Science, 5(3):261–278, October 2015. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-014-0101-x.

Brancazio:2023:MME

[BM23] Nick Brancazio and Russell Meyer. Minimal model explanations of cognition. European Journal for Philosophy of Science, 13(3):

??, September 2023. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-023-00547-4.

Bocchi:2022:BVP

[Boc22] Federica Bocchi. Biodiversity vs. paleodiversity measurements: the incommensurability problem. European Journal for Philosophy of Science, 12(4):??, December 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-022-00494-6.

Boge:2019:WCS

[Bog19] Florian J. Boge. Why computer simulations are not inferences, and in what sense they are experiments. European Journal for Philosophy of Science, 9(1):??, January 2019. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-018-0239-z. See cor-

rection [Bog22].

Boge:2021:IPI

[Bog21] Florian J. Boge. Incompatibility and the pessimistic induction: a challenge for selective realism. European Journal for Philosophy of Science, 11(2):??, June 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-021-00367-4.

Boge:2022:CWC

[Bog22] Florian J. Boge. Correction to: Why computer simulations are not inferences, and in what sense they are experiments. European Journal for Philosophy of Science, 12(4):??, December 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-022-00501-w. See [Bog19].

Bollhagen:2021:IER

[Bol21] Andrew Bollhagen. The inchworm episode: Reconstituting the phenomenon of kinesin motility. European Journal for Philosophy of Science, 11(2):??, June 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-021-00358-5.

Boon:2020:RDP

[Boo20] Mieke Boon. The role of disciplinary perspectives in an epistemology of scientific models. European Journal for Philosophy of Science, 10(3):??, October 2020. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-020-00295-9.

Borgerson:2011:ADC

[Bor11] Kirstin Borgerson. Amending and defending Critical Contextual Empiricism. European Journal for Philosophy of Science, 1(3): ??, October 2011. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-011-0035-5.

Boon:2022:EEI

[BOS22] Mieke Boon, Mariana Orozco, and Kishore Sivakumar. Epistemological and educational issues in teaching practice-oriented scientific research: roles for philosophers of science. European Journal for Philosophy of Science, 12(1):??, March 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-022-00447-z.

Bourrat:2019:VIM

[Bou19] Pierrick Bourrat. Variation of information as a measure of one-to-one causal specificity. European Journal for Philosophy of Science, 9(1):??, January 2019. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-018-0224-6.

Bourrat:2024:ACI

[Bou24] Pierrick Bourrat. Adding causality to the information-theoretic perspective on individuality. European Journal for Philosophy of Science, 14(1):??, March 2024. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-023-00566-1.

Brown:2015:CPI

[BP15] M. Bryson Brown and Graham Priest. Chunk and permeate II: Bohr’s hydrogen atom. European Journal for Philosophy of Science, 5(3):297–314, October 2015. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-014-0104-7.

Beni:2021:AFE

[BP21a] Majid D. Beni and Ahti-Veikko Pietarinen. Aligning the freeenergy principle with Peirce’s logic of science and economy of research. European Journal for Philosophy of Science, 11(3):??, September 2021. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-021-00408-y.

Bokulich:2021:DMR

[BP21b] Alisa Bokulich and Wendy Parker. Data models, representation and adequacy-for-purpose. European Journal for Philosophy of Science, 11(1):??, March 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-020-00345-2.

Brighouse:2014:GPA

[Bri14] Carolyn Brighouse. Geometric possibility- an argument from dimension. European Journal for Philosophy of Science, 4(1):31– 54, January 2014. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-013-0074-1.

Baras:2020:CEC

[BS20] Dan Baras and Orly Shenker. Calling for explanation: the case of the thermodynamic past state. European Journal for Philosophy of Science, 10(3):??, October 2020. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-020-00297-7.

Bursten:2021:BLT

[BS21] Julia R. S. Bursten and Matthew Strandmark. Better learning through history: using archival resources to teach healthcare ethics to science students. European Journal for Philosophy of Science, 11(3):??, September 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-021-00406-0.

Bschir:2020:PCE

[Bsc20] Karim Bschir. Perspectivism in current epigenetics. European Journal for Philosophy of Science, 10(3):??, October 2020. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-020-00302-z.

Boon:2019:EIR

[BV19] Mieke Boon and Sophie Van Baalen. Epistemology for interdisciplinary research — shifting philosophical paradigms of science. European Journal for Philosophy of Science, 9(1):??, January 2019. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-018-0242-4.

Barman:2021:IES

[BvE21] Kristian González Barman and Dingmar van Eck. IBE in engineering science — the case of malfunction explanation. European Journal for Philosophy of Science, 11(1):??, March 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-020-00325-6.

Cabrera:2021:SPT

[Cab21] Frank Cabrera. Second philosophy and testimonial reliability: Philosophy of science for STEM students. European Journal for Philosophy of Science, 11(3):??, September 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-021-00392-3.

Canali:2022:PAS

[Can22] Stefano Canali. A pragmatic approach to scientific change: transfer, alignment, influence. European Journal for Philosophy of Science, 12(3):??, September 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-022-00477-7.

Casini:2017:MT

[Cas17] Lorenzo Casini. Malfunctions and teleology. European Journal for Philosophy of Science, 7(2):319–335, May 2017. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-016-0163-z.

Crupi:2023:CPB

[CC23] Vincenzo Crupi and Fabrizio Calzavarini. Critique of pure Bayesian cognitive science: a view from the philosophy of science. European Journal for Philosophy of Science, 13(3):??, September 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-023-00533-w.

Cinti:2022:ER

[CCS22] Enrico Cinti, Alberto Corti, and Marco Sanchioni. On entanglement as a relation. European Journal for Philosophy of Science, 12 (1):??, March 2022. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-021-00439-5.

Cavalcanti:2023:CRF

[CDR23] Eric G. Cavalcanti, Andrea Di Biagio, and Carlo Rovelli. On the consistency of relative facts. European Journal for Philosophy of Science, 13(4):??, December 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-023-00551-8.

Cellucci:2018:DM

[Cel18] Carlo Cellucci. Definition in mathematics. European Journal for Philosophy of Science, 8(3):605–629, October 2018. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-018-0203-y.

Chen:2023:ECA

[CH23] Ruey-Lin Chen and Jonathon Hricko. Experimental criteria for accessing reality: Perrin’s experimental demonstration of atoms and molecules. European Journal for Philosophy of Science, 13 (1):??, March 2023. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-023-00517-w.

Chall:2020:MGS

[Cha20] Cristin Chall. Model-groups as scientific research programmes. European Journal for Philosophy of Science, 10(1):??, January 2020. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-019-0271-7.

Chung:2022:CYZ

[Chu22] Julianne Chung. Creativity and Yóu: the Zhuāngzǐ and scientific inquiry. European Journal for Philosophy of Science, 12(2): ??, June 2022. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-022-00448-y.

Currie:2016:MPS

[CK16] Adrian Currie and Anton Killin. Musical pluralism and the science of music. European Journal for Philosophy of Science, 6(1):9– 30, January 2016. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-015-0123-z.

Carrillo:2023:MPA

[CK23] Natalia Carrillo and Tarja Knuuttila. Mechanisms and the problem of abstract models. European Journal for Philosophy of Science, 13(3):??, September 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-023-00530-z.

Clatterbuck:2013:ETE

[Cla13] Hayley Clatterbuck. The epistemology of thought experiments: a non-eliminativist, non-platonic account. European Journal for Philosophy of Science, 3(3):309–329, October 2013. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-013-0069-y.

Colaco:2020:RSP

[Col20] David Colaço. Recharacterizing scientific phenomena. European Journal for Philosophy of Science, 10(2):??, May 2020. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-020-0279-z.

Conix:2018:ITO

[Con18] Stijn Conix. Integrative taxonomy and the operationalization of evolutionary independence. European Journal for Philosophy of Science, 8(3):587–603, October 2018. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-018-0202-z.

Contessa:2021:MIR

[Con21] Gabriele Contessa. On the mitigation of inductive risk. European Journal for Philosophy of Science, 11(3):??, September 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-021-00381-6.

Cartwright:2020:MLE

[CPW20] Nancy Cartwright, John Pemberton, and Sarah Wieten. Mechanisms, laws and explanation. European Journal for Philosophy of Science, 10(3):??, October 2020. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-020-00284-y. See replies [Kri20,

Kri21].

Crowther:2015:DER

[Cro15] Karen Crowther. Decoupling emergence and reduction in physics. European Journal for Philosophy of Science, 5(3):419–445, October 2015. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-015-0119-8.

Castellani:2023:NMA

[CS23] Elena Castellani and Giulia Schettino. Nested modalities in astrophysical modeling. European Journal for Philosophy of Science, 13 (1):??, March 2023. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-023-00511-2.

Corral-Villate:2020:SNI

[CV20] Amaia Corral-Villate. On Shackel’s nothing from infinity paradox. European Journal for Philosophy of Science, 10(3):??, October 2020. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-020-0277-1.

Danne:2021:DES

[Dan21] Nicholas Matthew Danne. A dialogue on the ethics of science: Henri Poincaré and Pope Francis. European Journal for Philosophy of Science, 11(3):??, September 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-021-00385-2.

Darrigol:2022:GME

[Dar22] Olivier Darrigol. Geometry, mechanics, and experience: a historico–philosophical musing. European Journal for Philosophy of Science, 12(4):??, December 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-022-00491-9.

DeHaro:2018:ITS

[DdR18] Sebastian De Haro and Henk W. de Regt. Interpreting theories without a spacetime. European Journal for Philosophy of Science, 8(3):631–670, October 2018. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-018-0204-x.

DeHaro:2019:TTE

[De 19] Sebastian De Haro. Towards a theory of emergence for the physical sciences. European Journal for Philosophy of Science, 9(3): ??, October 2019. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-019-0261-9.

DeBenedetto:2021:ETS

[De 21] Matteo De Benedetto. Explication as a Three-Step Procedure: the case of the Church–Turing Thesis. European Journal for Philosophy of Science, 11(1):??, March 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-020-00337-2.

DeBenedetto:2023:LEP

[De 23] Matteo De Benedetto. Lakatosian and Euclidean populations: a pluralist approach to conceptual change in mathematics. European Journal for Philosophy of Science, 13(3):??, September 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-023-00543-8.

Dethier:2024:CCA

[Det24] Corey Dethier. Contrast classes and agreement in climate modeling. European Journal for Philosophy of Science, 14(1):??, March 2024. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-024-00577-6.

DelSanto:2023:PRR

[DG23] Flavio Del Santo and Nicolas Gisin. Potentiality realism: a realistic and indeterministic physics based on propensities. European Journal for Philosophy of Science, 13(4):??, December 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-023-00561-6.

Dorato:2011:E

[DH11] Mauro Dorato and Carl Hoefer. Editorial. European Journal for Philosophy of Science, 1(1):1–2, January 2011. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-011-0021-y.

Diez:2018:FTB

[Díe18] José Díez. A (Fatal) Trilemma for best theory realism. European Journal for Philosophy of Science, 8(2):271–291, May 2018. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-017-0185-1.

DiFrisco:2018:TPF

[DiF18] James DiFrisco. Token physicalism and functional individuation. European Journal for Philosophy of Science, 8(3):309–329, October 2018. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-017-0188-y.

Dimitrakos:2020:RRR

[Dim20a] Thodoris Dimitrakos. Reconstructing rational reconstructions: on Lakatos’s account on the relation between history and philosophy of science. European Journal for Philosophy of Science, 10(3): ??, October 2020. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-020-00293-x.

Dimitrijevic:2020:CCP

[Dim20b] Dejan R. Dimitrijević. Causal closure of the physical, mental causation, and physics. European Journal for Philosophy of Science, 10(1):??, January 2020. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-019-0267-3.

Dupre:2022:PEC

[DL22] John Dupré and Sabina Leonelli. Process epistemology in the COVID-19 era: rethinking the research process to avoid dangerous forms of reification. European Journal for Philosophy of Science, 12(1):??, March 2022. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-022-00450-4.

Dawid:2023:TVI

[DM23] Richard Dawid and Casey McCoy. Testability and viability: is inflationary cosmology “Scientific”? European Journal for Philosophy of Science, 13(4):??, December 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-023-00556-3.

deMelo-Martin:2023:SRS

[dMMI23] Inmaculada de Melo-Martín and Kristen Intemann. Socially responsible science: Exploring the complexities. European Journal for Philosophy of Science, 13(3):??, September 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-023-00537-6.

deOliveira:2022:RA

[dO22] Guilherme Sanches de Oliveira. Radical artifactualism. European Journal for Philosophy of Science, 12(2):??, June 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-022-00462-0.

Doboszewski:2019:RSD

[Dob19] Juliusz Doboszewski. Relativistic spacetimes and definitions of determinism. European Journal for Philosophy of Science, 9(2):??, May 2019. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-019-0248-6.

Dolev:2018:PST

[Dol18] Yuval Dolev. Physics’ silence on time. European Journal for Philosophy of Science, 8(3):455–469, October 2018. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-017-0195-z.

Doppelt:2014:BTS

[Dop14] Gerald Doppelt. Best theory scientific realism. European Journal for Philosophy of Science, 4(2):271–291, May 2014. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-014-0090-9.

Douven:2023:PFL

[Dou23] Igor Douven. Pandemics and flexible lockdowns: In praise of agent-based modeling. European Journal for Philosophy of Science, 13(3):??, September 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-023-00541-w.

Dowe:2019:ATD

[Dow19] Phil Dowe. Action at a temporal distance in the best systems account. European Journal for Philosophy of Science, 9(3):??, October 2019. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-019-0258-4.

Dragulinescu:2017:IBE

[Dra17] Stefan Dragulinescu. Inference to the best explanation as a theory for the quality of mechanistic evidence in medicine. European Journal for Philosophy of Science, 7(2):353–372, May 2017. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-016-0165-x.

deRegt:2021:HWP

[dRK21] Henk W. de Regt and Edwin Koster. The hows and whys of philosophy of science teaching: a comparative analysis. European Journal for Philosophy of Science, 11(4):??, December 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-021-00421-1.

Diez:2023:HDN

[DS23] José Díez and Javier Suárez. How do networks explain? A neo-Hempelian approach to network explanations of the ecology of the microbiome. European Journal for Philosophy of Science, 13(3): ??, September 2023. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-023-00549-2.

Due:2023:WSE

[Due23] Austin Due. What are side effects? European Journal for Philosophy of Science, 13(1):??, March 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-023-00519-8.

Dunleavy:2022:PDJ

[Dun22] Daniel J. Dunleavy. Progressive and degenerative journals: on the growth and appraisal of knowledge in scholarly publishing. European Journal for Philosophy of Science, 12(4):??, December 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-022-00492-8.

Dupre:2011:ESN

[Dup11] John Dupré. Emerging sciences and new conceptions of disease; or, beyond the monogenomic differentiated cell lineage. European Journal for Philosophy of Science, 1(1):119–131, January 2011. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-010-0008-0.

Egg:2014:NLC

[EE14] Matthias Egg and Michael Esfeld. Non-local common cause explanations for EPR. European Journal for Philosophy of Science, 4(2): 181–196, May 2014. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-013-0081-2.

Elliott-Graves:2022:WGM

[EG22a] Alkistis Elliott-Graves. What are general models about? European Journal for Philosophy of Science, 12(4):??, December 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-022-00502-9.

Everth:2022:ERD

[EG22b] Thomas Everth and Laura Gurney. Emergent realities: Diffracting Barad within a quantum-realist ontology of matter and politics. European Journal for Philosophy of Science, 12(3):??, September 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-022-00476-8.

Egg:2021:QOS

[Egg21] Matthias Egg. Quantum ontology without speculation. European Journal for Philosophy of Science, 11(1):??, March 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-020-00346-1.

ElMawas:2021:POC

[El 21] Omar El Mawas. Probing ‘operational coherence’ in Hasok Chang’s pragmatic realism. European Journal for Philosophy of Science, 11(4):??, December 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-021-00425-x.

Eronen:2012:PPC

[Ero12] Markus I. Eronen. Pluralistic physicalism and the causal exclusion argument. European Journal for Philosophy of Science, 2(2):219– 232, May 2012. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-011-0041-7.

Esfeld:2013:OSR

[Esf13] Michael Esfeld. Ontic structural realism and the interpretation of quantum mechanics. European Journal for Philosophy of Science, 3(1):19–32, January 2013. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-012-0054-x.

Evans:2020:WCB

[ET20] Peter W. Evans and Karim P. Y. Thébault. What can bouncing oil droplets tell us about quantum mechanics? European Journal for Philosophy of Science, 10(3):??, October 2020. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-020-00301-0.

Eva:2016:CTP

[Eva16] Benjamin Eva. Category theory and physical structuralism. European Journal for Philosophy of Science, 6(2):231–246, May 2016. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-015-0129-6.

Evans:2020:PO

[Eva20] Peter W. Evans. Perspectival objectivity. European Journal for Philosophy of Science, 10(2):??, May 2020. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-020-00286-w.

Eigi-Watkin:2022:ANE

[EW22] Jaana Eigi-Watkin. Applying the notion of epistemic risk to argumentation in philosophy of science. European Journal for Philosophy of Science, 12(2):??, June 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-022-00455-z.

Eigi-Watkin:2024:HIR

[EWVTL24] Jaana Eigi-Watkin, Katrin Velbaum, Edit Talpsepp, and Endla Lõhkivi. How interdisciplinary researchers see themselves: plurality of understandings of interdisciplinarity within a field and why it matters. European Journal for Philosophy of Science, 14 (1):??, March 2024. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-024-00572-x.

Fagerberg:2022:RNK

[Fag22] Harriet Fagerberg. Reactive natural kinds and varieties of dependence. European Journal for Philosophy of Science, 12(4):??, December 2022. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-022-00500-x.

Farr:2020:ETQ

[Far20] Matt Farr. Explaining temporal qualia. European Journal for Philosophy of Science, 10(1):??, January 2020. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-019-0264-6.

Fatollahi:2023:PMS

[Fat23] Alireza Fatollahi. Predictivism and model selection. European Journal for Philosophy of Science, 13(1):??, March 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-023-00512-1.

Field:2021:EAE

[FD21] Sarahanne M. Field and Maarten Derksen. Experimenter as automaton; experimenter as human: exploring the position of the researcher in scientific research. European Journal for Philosophy of Science, 11(1):??, March 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-020-00324-7.

Feest:2020:CVP

[Fee20] Uljana Feest. Construct validity in psychological tests — the case of implicit social cognition. European Journal for Philosophy of Science, 10(1):??, January 2020. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-019-0270-8.

Feest:2022:DQE

[Fee22] Uljana Feest. Data quality, experimental artifacts, and the reactivity of the psychological subject matter. European Journal for Philosophy of Science, 12(1):??, March 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-021-00443-9.

Feldbacher-Escamilla:2019:MCA

[FEG19] Christian J. Feldbacher-Escamilla and Alexander Gebharter. Modeling creative abduction Bayesian style. European Journal for Philosophy of Science, 9(1):??, January 2019. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-018-0234-4.

Fellowes:2021:SPP

[Fel21] Sam Fellowes. Scientific Perspectivism and psychiatric diagnoses: respecting history and constraining relativism. European Journal for Philosophy of Science, 11(1):??, March 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-020-00320-x.

Field:2022:SQT

[Fie22] Grace E. Field. On the status of quantum tunnelling time. European Journal for Philosophy of Science, 12(4):??, December 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-022-00483-9.

Fillion:2021:SLS

[Fil21] Nicolas Fillion. Semantic layering and the success of mathematical sciences. European Journal for Philosophy of Science, 11(3):??, September 2021. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-021-00394-1.

Finn:2021:MMP

[Fin21] Suki Finn. Methodology for the metaphysics of pregnancy. European Journal for Philosophy of Science, 11(3):??, September 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-021-00378-1.

Fischer:2021:PPI

[Fis21] Marc Fischer. On the principal principle and imprecise subjective Bayesianism. European Journal for Philosophy of Science, 11(2): ??, June 2021. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-021-00356-7.

Fischer:2024:BBD

[Fis24] Enno Fischer. Broken brakes and dreaming drivers: the heuristic value of causal models in the law. European Journal for Philosophy of Science, 14(1):??, March 2024. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-023-00563-4.

Fletcher:2012:WCN

[Fle12] Samuel Craig Fletcher. What counts as a Newtonian system? The view from Norton’s dome. European Journal for Philosophy of Science, 2(3):275–297, October 2012. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-011-0040-8.

Fletcher:2019:SRE

[Fle19] Samuel C. Fletcher. Stopping rules as experimental design. European Journal for Philosophy of Science, 9(2):??, May 2019. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-019-0252-x.

Fletcher:2021:HMR

[Fle21a] Samuel C. Fletcher. How (not) to measure replication. European Journal for Philosophy of Science, 11(2):??, June 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-021-00377-2.

Fletcher:2021:RRP

[Fle21b] Samuel C. Fletcher. The role of replication in psychological science. European Journal for Philosophy of Science, 11(1):??, March 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-020-00329-2.

Fillion:2018:EAB

[FM18] Nicolas Fillion and Robert H. C. Moir. Explanation and abstraction from a backward-error analytic perspective. European Journal for Philosophy of Science, 8(3):735–759, October 2018. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-018-0208-6.

Frigg:2018:TVR

[FN18] Roman Frigg and James Nguyen. The turn of the valve: representing with material models. European Journal for Philosophy of Science, 8(2):205–224, May 2018. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-017-0182-4.

Forgione:2024:GFT

[For24] Marco Forgione. Group field theories: decoupling spacetime emergence from the ontology of non-spatiotemporal entities. European Journal for Philosophy of Science, 14(2):??, June 2024. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-024-00585-6.

French:2020:ISP

[Fre20] Steven French. Imagination in Scientific Practice. European Journal for Philosophy of Science, 10(3):??, October 2020. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-020-00291-z.

Freeborn:2023:EFT

[Fre23] David Freeborn. Efficiency and fairness trade-offs in two player bargaining games. European Journal for Philosophy of Science, 13(4):??, December 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-023-00553-6.

Frisch:2012:NPC

[Fri12] Mathias Frisch. No place for causes? Causal skepticism in physics. European Journal for Philosophy of Science, 2(3):313–336, October 2012. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-011-0044-4.

Friederich:2013:GSB

[Fri13] Simon Friederich. Gauge symmetry breaking in gauge theories — in search of clarification. European Journal for Philosophy of Science, 3(2):157–182, May 2013. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-012-0061-y.

Frisch:2015:POE

[Fri15] Mathias Frisch. Predictivism and old evidence: a critical look at climate model tuning. European Journal for Philosophy of Science, 5(2):171–190, May 2015. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-015-0110-4.

Friend:2016:LC

[Fri16] Toby Friend. Laws are conditionals. European Journal for Philosophy of Science, 6(1):123–144, January 2016. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-015-0131-z.

Friend:2022:HPT

[Fri22] Toby Friend. The Humean pragmatic turn and the case for revisionary best systems accounts. European Journal for Philosophy of Science, 12(1):??, March 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-021-00436-8.

Fernandez-Roldan:2024:ESR

[FRT24] Alejandro Fernández-Roldan and David Teira. The epistemic status of reproducibility in political fact-checking. European Journal for Philosophy of Science, 14(1):??, March 2024. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-024-00575-8.

Fry:2016:BEU

[Fry16] Richard J. Fry. Backwards explanation and unification. European Journal for Philosophy of Science, 6(1):55–70, January 2016. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-015-0121-1.

Fuller:2021:IPD

[FS21] Gareth P. Fuller and Armin W. Schulz. Idealizations and partitions: a defense of robustness analysis. European Journal for Philosophy of Science, 11(4):??, December 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-021-00428-8.

Frigg:2022:IAH

[FS22] Roman Frigg and Leonard A. Smith. An ineffective antidote for hawkmoths. European Journal for Philosophy of Science, 12(2): ??, June 2022. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-022-00459-9.

Fumagalli:2016:CMR

[Fum16] Roberto Fumagalli. Choice models and realistic ontologies: three challenges to neuro-psychological modellers. European Journal for Philosophy of Science, 6(1):145–164, January 2016. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-015-0134-9.

Frigg:2011:EYA

[FV11] Roman Frigg and Ioannis Votsis. Everything you always wanted to know about structural realism but were afraid to ask. European Journal for Philosophy of Science, 1(2):227–276, May 2011. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-011-0025-7.

Fedyk:2021:CPV

[FX21] Mark Fedyk and Fei Xu. Creativity as potentially valuable improbable constructions. European Journal for Philosophy of Science, 11 (1):??, March 2021. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-020-00343-4.

Green:2021:APB

[GAD^{+}21] Sara Green, Hanne Andersen, Kristian Danielsen, Claus Emmeche, Christian Joas, Mikkel Willum Johansen, Caio Nagayoshi, Joeri Witteveen, and Henrik Kragh Sørensen. Adapting practicebased philosophy of science to teaching of science students. European Journal for Philosophy of Science, 11(3):??, September 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-021-00393-2.

Gomori:2023:CBT

[GH23] Márton Gömöri and Carl Hoefer. Classicality and Bell’s theorem. European Journal for Philosophy of Science, 13(3):??, September 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-023-00531-y.

Ghomi:2023:IMS

[Gho23] Hamed Tabatabaei Ghomi. Irrational methods suggest indecomposability and emergence. European Journal for Philosophy of Science, 13(1):??, March 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-022-00503-8.

Giovanelli:2018:PKM

[Gio18] Marco Giovanelli. ‘Physics is a kind of metaphysics’: Émile Meyerson and Einstein’s late rationalistic realism. European Journal for Philosophy of Science, 8(3):783–829, October 2018. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-018-0211-y.

Giovanelli:2021:NCP

[Gio21] Marco Giovanelli. Nothing but coincidences: the point-coincidence and Einstein’s struggle with the meaning of coordinates in physics. European Journal for Philosophy of Science, 11(2):??, June 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/

s13194-020-00332-7. See correction [Gio22].

Giovanelli:2022:CNC

[Gio22] Marco Giovanelli. Correction to: Nothing but coincidences: the point-coincidence and Einstein’s struggle with the meaning of coordinates in physics. European Journal for Philosophy of Science, 12(3):??, September 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-022-00482-w. See [Gio21].

Giovanelli:2023:ARE

[Gio23] Marco Giovanelli. Appearance and reality: Einstein and the early debate on the reality of length contraction. European Journal for Philosophy of Science, 13(4):??, December 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-023-00555-4.

Gagne-Julien:2023:SSP

[GJB23] Anne-Marie Gagné-Julien and Andréanne Bérubé. Should social pragmatic communication disorder be included in DSM-5? On uncertainties, pragmatic considerations, and the psychiatric kind debate. European Journal for Philosophy of Science, 13(3):??, September 2023. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-023-00535-8.

Giorgi:2021:SCP

[GL21] Rodolfo Giorgi and Andrea Lavazza. The self and its causal powers between metaphysics and science. European Journal for Philosophy of Science, 11(1):??, March 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-020-00316-7.

Glick:2024:MIE

[GL24] David Glick and Baptiste Le Bihan. Metaphysical indeterminacy in Everettian quantum mechanics. European Journal for Philosophy of Science, 14(1):??, March 2024. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-023-00562-5.

Green:2015:DSA

[GLB15] Sara Green, Arnon Levy, and William Bechtel. Design sans adaptation. European Journal for Philosophy of Science, 5(1):15–29, January 2015. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-014-0096-3.

Glick:2021:QLS

[Gli21] David Glick. QBism and the limits of scientific realism. European Journal for Philosophy of Science, 11(2):??, June 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-021-00366-5.

Godman:2022:WSS

[GM22] Marion Godman and Caterina Marchionni. What should scientists do about (harmful) interactive effects? European Journal for Philosophy of Science, 12(4):??, December 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-022-00493-7.

Gomes:2021:HES

[Gom21] Henrique Gomes. Holism as the empirical significance of symmetries. European Journal for Philosophy of Science, 11(3):??, September 2021. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-021-00397-y.

Garcia-Portela:2023:OEA

[GPM23] Laura García-Portela and Douglas Maraun. Overstating the effects of anthropogenic climate change? A critical assessment of attribution methods in climate science. European Journal for Philosophy of Science, 13(1):??, March 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-023-00516-x.

Grantham:2024:TMN

[Gra24] Todd Grantham. Toward a more natural historical attitude. European Journal for Philosophy of Science, 14(1):??, March 2024. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-023-00564-3.

Gillies:2013:SCM

[GS13] Donald Gillies and Aidan Sudbury. Should causal models always be Markovian? The case of multi-causal forks in medicine. European Journal for Philosophy of Science, 3(3):275–308, October 2013. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-013-0068-z.

Guay:2016:NLE

[GS16] Alexandre Guay and Olivier Sartenaer. A new look at emergence. Or when after is different. European Journal for Philosophy of Science, 6(2):297–322, May 2016. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-016-0140-6.

Gardenfors:2018:IKW

[GS18]  Peter Gärdenfors and Andreas Stephens.  Induction and

knowledge-what. European Journal for Philosophy of Science, 8(3):471–491, October 2018. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-017-0196-y.

Gefaell:2022:CIE

[GS22a] Juan Gefaell and Cristian Saborido. Correction to: Incommensurability and the Extended Evolutionary Synthesis: taking Kuhn seriously. European Journal for Philosophy of Science, 12(2):??, June 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-022-00460-2. See [GS22b].

Gefaell:2022:IEE

[GS22b] Juan Gefaell and Cristian Saborido. Incommensurability and the Extended Evolutionary Synthesis: taking Kuhn seriously. European Journal for Philosophy of Science, 12(2):??, June 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/

s13194-022-00456-y. See correction [GS22a].

Guala:2023:SKH

[Gua23] Francesco Guala. Social kinds: historical and multi-functional. European Journal for Philosophy of Science, 13(3):??, September 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-023-00534-9.

Guralp:2020:EAU

[Gur20] Genco Guralp. The evidence for the accelerating universe: endorsement and robust consistency. European Journal for Philosophy of Science, 10(2):??, May 2020. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-020-0276-2.

Guttinger:2020:LR

[Gut20] Stephan Guttinger. The limits of replicability. European Journal for Philosophy of Science, 10(2):??, May 2020. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-019-0269-1.

Grune-Yanoff:2014:TPS

[GY14] Till Grüne-Yanoff. Teaching philosophy of science to scientists: why, what and how. European Journal for Philosophy of Science, 4(1):115–134, January 2014. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-013-0078-x.

Grune-Yanoff:2016:ISI

[GY16] Till Grüne-Yanoff. Interdisciplinary success without integration. European Journal for Philosophy of Science, 6(3):343–360, October 2016. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-016-0139-z.

Grune-Yanoff:2022:OFC

[GY22] Till Grüne-Yanoff. Overcoming Frege’s curse: heuristic reasoning as the basis for teaching philosophy of science to scientists. European Journal for Philosophy of Science, 12(1):??, March 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-022-00452-2.

Hacohen:2022:PAH

[Hac22] Ori Hacohen. The problem with appealing to history in defining neural representations. European Journal for Philosophy of Science, 12(3):??, September 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-022-00473-x.

Haggqvist:2022:NWS

[Häg22] Sören Häggqvist. No, water (still) doesn’t have a microstructural essence (reply to Hoefer & Martí). European Journal for Philosophy of Science, 12(2):??, June 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-022-00461-1. See [HM19].

Hands:2016:DRC

[Han16] D. Wade Hands. Derivational robustness, credible substitute systems and mathematical economic models: the case of stability analysis in Walrasian general equilibrium theory. European Journal for Philosophy of Science, 6(1):31–53, January 2016. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-015-0130-0.

Hansson:2019:FES

[Han19] Sven Ove Hansson. Farmers’ experiments and scientific methodology. European Journal for Philosophy of Science, 9(3):??, October 2019. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-019-0255-7.

Hansson:2020:SCC

[Han20] Sven Ove Hansson. Social constructionism and climate science denial. European Journal for Philosophy of Science, 10(3):??, October 2020. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-020-00305-w. See reply [VNT22].

Hansson:2022:MAA

[Han22] Sven Ove Hansson. Misconstrued arguments about cultural theory. European Journal for Philosophy of Science, 12(3):??, September 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-022-00472-y.

Han:2023:TMP

[Han23] HyeJeong Han. Taking model pursuit seriously. European Journal for Philosophy of Science, 13(2):??, June 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-023-00524-x.

Harbecke:2019:TCB

[Har19] Jens Harbecke. Two challenges for a boolean approach to constitutive inference. European Journal for Philosophy of Science, 9(1): ??, January 2019. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-018-0238-0.

Heaphy:2015:RCM

[Hea15] Liam James Heaphy. The role of climate models in adaptation decision-making: the case of the UK climate projections 2009. European Journal for Philosophy of Science, 5(2):233–257, May 2015. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-015-0114-0.

Heathcote:2021:CSI

[Hea21] Adrian Heathcote. Countability and self-identity. European Journal for Philosophy of Science, 11(4):??, December 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-021-00423-z.

Heilmann:2014:SCN

[Hei14] Conrad Heilmann. Success conditions for nudges: a methodological critique of libertarian paternalism. European Journal for Philosophy of Science, 4(1):75–94, January 2014. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-013-0076-z.

Helgeson:2017:POD

[Hel17] Casey Helgeson. Pattern as observation: Darwin’s ‘great facts’ of geographical distribution. European Journal for Philosophy of Science, 7(2):337–351, May 2017. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-016-0164-y.

Henschen:2018:PIC

[Hen18] Tobias Henschen. The in-principle inconclusiveness of causal evidence in macroeconomics. European Journal for Philosophy of Science, 8(3):709–733, October 2018. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-018-0207-7.

Hensel:2020:DTC

[Hen20] Witold M. Hensel. Double trouble? The communication dimension of the reproducibility crisis in experimental psychology and neuroscience. European Journal for Philosophy of Science, 10(3): ??, October 2020. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-020-00317-6.

Henschen:2021:HSA

[Hen21] Tobias Henschen. How strong is the argument from inductive risk? European Journal for Philosophy of Science, 11(3):??, September 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-021-00409-x.

Herzberg:2014:NNA

[Her14] Frederik Herzberg. A note on “The no alternatives argument” by Richard Dawid, Stephan Hartmann and Jan Sprenger. European Journal for Philosophy of Science, 4(3):375–384, October 2014. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-014-0092-7.

Hoyningen-Huene:2023:OVF

[HH23] Paul Hoyningen-Huene. Objectivity, value-free science, and inductive risk. European Journal for Philosophy of Science, 13(1): ??, March 2023. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-023-00518-9.

Hijmans:2022:ACE

[Hij22] Sarah N. Hijmans. Analogy and composition in Early Nineteenth-Century chemistry the case of aluminium. European Journal for Philosophy of Science, 12(1):??, March 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-021-00442-w.

Hilligardt:2022:LBV

[Hil22] Hannah Hilligardt. Looking beyond values: The legitimacy of social perspectives, opinions and interests in science. European Journal for Philosophy of Science, 12(4):??, December 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-022-00490-w.

Hireche:2021:VDE

[Hir21] Salim Hirèche. Varieties of dispositional essentialism about natural laws. European Journal for Philosophy of Science, 11(3):??, September 2021. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-021-00375-4.

Hitchcock:2012:TER

[Hit12] Christopher Hitchcock. Thought experiments, real experiments, and the expertise objection. European Journal for Philosophy of Science, 2(2):205–218, May 2012. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-012-0051-0.

Hireche:2021:MSL

[HLMV21] Salim Hirèche, Niels Linnemann, Robert Michels, and Lisa Vogt. The modal status of the laws of nature. Tahko’s hybrid view and the kinematical/dynamical distinction. European Journal for Philosophy of Science, 11(1):??, March 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-020-00335-4.

Hoefer:2019:WME

[HM19] Carl Hoefer and Genoveva Martí. Water has a microstructural essence after all. European Journal for Philosophy of Science, 9 (1):??, January 2019. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-018-0236-2. See reply [Häg22].

Hoefer:2020:RRP

[HM20] Carl Hoefer and Genoveva Martí. Realism, reference & perspective. European Journal for Philosophy of Science, 10(3):??, October 2020. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-020-00306-9.

Huggett:2021:LHM

[HM21] Nick Huggett and Keizo Matsubara. Lost horizon? — Modeling black holes in string theory. European Journal for Philosophy of Science, 11(3):??, September 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-021-00376-3.

Hochstein:2022:FBN

[Hoc22] Eric Hochstein. Foregrounding and backgrounding: a new interpretation of “levels” in science. European Journal for Philosophy of Science, 12(2):??, June 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-022-00457-x.

Holtgen:2021:SST

[Höl21] Benedikt T. A. Höltgen. Structure-sensitive testimonial norms. European Journal for Philosophy of Science, 11(3):??, September 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-021-00360-x.

Howson:2017:RIT

[How17] Colin Howson. Regularity and infinitely tossed coins. European Journal for Philosophy of Science, 7(1):97–102, January 2017. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-016-0147-z.

Hubert:2018:WFM

[HR18] Mario Hubert and Davide Romano. The wave-function as a multifield. European Journal for Philosophy of Science, 8(3):521–537, October 2018. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-017-0198-9.

Hartmann:2012:FPS

[HS12] Stephan Hartmann and Jan Sprenger. The future of philosophy of science: introduction. European Journal for Philosophy of Science, 2(2):157–159, May 2012. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-012-0052-z.

Harbecke:2019:REC

[HS19a] Jens Harbecke and Oron Shagrir. The role of the environment in computational explanations. European Journal for Philosophy of Science, 9(3):??, October 2019. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-019-0263-7.

Hatleback:2019:RGM

[HS19b] Eric Nelson Hatleback and Jonathan M. Spring. A refinement to the general mechanistic account. European Journal for Philosophy of Science, 9(2):??, May 2019. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-018-0237-1.

Hofer-Szabo:2021:TNH

[HS21] Gábor Hofer-Szabó. Three noncontextual hidden variable models for the Peres–Mermin square. European Journal for Philosophy of Science, 11(1):??, March 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-020-00339-0.

Henne:2024:PVS

[HTS24] Céline Henne, Hannah Tomczyk, and Christoph Sperber. Physicists’ views on scientific realism. European Journal for Philosophy of Science, 14(1):??, March 2024. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-024-00570-z.

Huber:2017:JDI

[Hub17] Franz Huber. On the justification of deduction and induction. European Journal for Philosophy of Science, 7(3):507–534, October 2017. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-017-0177-1.

Hubert:2021:UPW

[Hub21] Mario Hubert. Understanding physics: ‘What?’, ‘Why?’, and ‘How?’. European Journal for Philosophy of Science, 11(3):??, September 2021. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-021-00399-w.

Huebner:2011:GCE

[Hue11] Bryce Huebner. Genuinely collective emotions. European Journal for Philosophy of Science, 1(1):89–118, January 2011. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-010-0006-2.

Hocquet:2021:EIC

[HW21] Alexandre Hocquet and Frédéric Wieber. Epistemic issues in computational reproducibility: software as the elephant in the room. European Journal for Philosophy of Science, 11(2):??, June 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-021-00362-9.

Intemann:2014:APP

[IdMM14] Kristen Intemann and Inmaculada de Melo-Martín. Addressing problems in profit-driven research: how can feminist conceptions of objectivity help? European Journal for Philosophy of Science, 4(2):135–151, May 2014. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-013-0079-9.

Ijas:2021:EBP

[IK21] Tero Ijäs and Rami Koskinen. Exploring biological possibility through synthetic biology. European Journal for Philosophy of Science, 11(2):??, June 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-021-00364-7.

Ioannidis:2021:NLT

[ILP21] Stavros Ioannidis, Vassilis Livanios, and Stathis Psillos. No laws and (thin) powers in, no (governing) laws out. European Journal for Philosophy of Science, 11(1):??, March 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-020-00304-x.

Ivani:2022:PEA

[IN22] Silvia Ivani and Catarina Dutilh Novaes. Public engagement and argumentation in science. European Journal for Philosophy of Science, 12(3):??, September 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-022-00480-y.

Intemann:2015:DBL

[Int15] Kristen Intemann. Distinguishing between legitimate and illegitimate values in climate modeling. European Journal for Philosophy of Science, 5(2):217–232, May 2015. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-014-0105-6.

Illari:2017:EL

[IR17] Phyllis Illari and Federica Russo. Editors’ letter. European Journal for Philosophy of Science, 7(3):391–392, October 2017. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-017-0186-0.

Illari:2018:ELa

[IR18a] Phyllis Kirstin Illari and Federica Russo. Editors’ letter. European Journal for Philosophy of Science, 8(1):1–2, January 2018. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-017-0199-8.

Illari:2018:ELb

[IR18b] Phyllis Kirstin Illari and Federica Russo. Editors’ letter. European Journal for Philosophy of Science, 8(3):307–308, October 2018. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-018-0220-x.

Ivani:2019:WWS

[Iva19] Silvia Ivani. What we (should) talk about when we talk about fruitfulness. European Journal for Philosophy of Science, 9(1): ??, January 2019. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-018-0231-7.

Illari:2012:WMT

[IW12] Phyllis McKay Illari and Jon Williamson. What is a mechanism? Thinking about mechanisms across the sciences. European Journal for Philosophy of Science, 2(1):119–135, January 2012. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-011-0038-2.

Jacoby:2022:EMI

[Jac22] Franklin R. Jacoby. Exploratory modeling and indeterminacy in the search for life. European Journal for Philosophy of Science, 12 (2):??, June 2022. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-022-00469-7.

Jaksland:2020:OPN

[Jak20] Rasmus Jaksland. Old problems for neo-positivist naturalized metaphysics. European Journal for Philosophy of Science, 10(2): ??, May 2020. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-020-00282-0.

Jaksland:2021:ACB

[Jak21a] Rasmus Jaksland. An apology for conflicts between metaphysics and science in naturalized metaphysics. European Journal for Philosophy of Science, 11(3):??, September 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-021-00390-5.

Jaksland:2021:TSC

[Jak21b] Rasmus Jaksland. Teaching scientific creativity through philosophy of science. European Journal for Philosophy of Science, 11(4): ??, December 2021. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-021-00427-9.

Jaksland:2023:DAR

[Jak23a] Rasmus Jaksland. Decoherence, appearance, and reality in agential realism. European Journal for Philosophy of Science, 13(2): ??, June 2023. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-023-00528-7.

Jaksland:2023:DTU

[Jak23b] Rasmus Jaksland. Distinguishing two (unsound) arguments for quantum social science. European Journal for Philosophy of Science, 13(3):??, September 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-023-00540-x.

Jebeile:2021:MSP

[JB21a] Julie Jebeile and Anouk Barberousse. Model spread and progress in climate modelling. European Journal for Philosophy of Science, 11(3):??, September 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-021-00387-0.

Jimenez-Buedo:2021:RSS

[JB21b] María Jiménez-Buedo. Reactivity in social scientific experiments: what is it and how is it different (and worse) than a Placebo effect? European Journal for Philosophy of Science, 11(2):??, June 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-021-00350-z.

Joaquim:2019:PBP

[JFEH19] Leyla Mariane Joaquim, Olival Freire, Jr., and Charbel N. El-Hani. From physics to biology: physicists in the search for systemic biological explanations. European Journal for Philosophy of Science, 9(2):??, May 2019. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-019-0254-8.

Johnson:2012:RBP

[Joh12] Gregory Johnson. The relationship between psychological capacities and neurobiological activities. European Journal for Philosophy of Science, 2(3):453–480, October 2012. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-012-0053-y.

John:2015:EEI

[Joh15a] Stephen John. Erratum to: The example of the IPCC does not vindicate the Value Free Ideal: a reply to Gregor Betz. European Journal for Philosophy of Science, 5(2):259, May 2015. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-015-0106-0. See [Joh15b].

John:2015:EID

[Joh15b] Stephen John. The example of the IPCC does not vindicate the Value Free Ideal: a reply to Gregor Betz. European Journal for Philosophy of Science, 5(1):1–13, January 2015. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/

s13194-014-0095-4. See erratum [Joh15a].

Jaffro:2021:UC

[JP21] Eva Jaffro and Cédric Paternotte. Unexplained cooperation. European Journal for Philosophy of Science, 11(3):??, September 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-021-00386-1.

Justus:2012:CCD

[Jus12] James Justus. Carnap on concept determination: methodology for philosophy of science. European Journal for Philosophy of Science, 2(2):161–179, May 2012. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-011-0027-5.

Karaca:2021:VIR

[Kar21] Koray Karaca. Values and inductive risk in machine learning modelling: the case of binary classification models. European Journal for Philosophy of Science, 11(4):??, December 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-021-00405-1.

Kashyap:2023:GRM

[Kas23] Abhishek Kashyap. General relativity, MOND, and the problem of unconceived alternatives. European Journal for Philosophy of Science, 13(3):??, September 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-023-00532-x.

Katzav:2023:EPC

[Kat23] Joel Katzav. Epistemic possibilities in climate science: lessons from some recent research in the context of discovery. European Journal for Philosophy of Science, 13(4):??, December 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-023-00560-7.

Knuuttila:2011:HDM

[KB11] Tarja Knuuttila and Mieke Boon. How do models give us knowledge? The case of Carnot’s ideal heat engine. European Journal for Philosophy of Science, 1(3):??, October 2011. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-011-0029-3.

Kuhlmann:2014:RBQ

[KG14] Meinard Kuhlmann and Stuart Glennan. On the relation between quantum mechanical and neo-mechanistic ontologies and explanatory strategies. European Journal for Philosophy of Science, 4(3):337–359, October 2014. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-014-0088-3.

Keyser:2021:MPP

[KH21] Vadim Keyser and Hannah Howland. Measurement perspective, process, and the pandemic. European Journal for Philosophy of Science, 11(1):??, March 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-020-00326-5.

Khalifa:2015:EDR

[Kha15] Kareem Khalifa. EMU defended: reply to Newman (2014). European Journal for Philosophy of Science, 5(3):377–385, October 2015. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-015-0112-2. See [New14].

Khalili:2022:PHA

[Kha22] Mahdi Khalili. From phenomenological-hermeneutical approaches to realist perspectivism. European Journal for Philosophy of Science, 12(4):??, December 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-022-00495-5.

Khosrowi:2023:EEC

[Kho23] Donal Khosrowi. Extrapolating from experiments, confidently. European Journal for Philosophy of Science, 13(2):??, June 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-023-00520-1.

Kidd:2021:CSA

[Kid21] Ian James Kidd. Creativity in science and the ‘anthropological turn’ in virtue theory. European Journal for Philosophy of Science, 11(1):??, March 2021. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-020-00334-5.

King:2020:ECE

[Kin20] Martin King. Explanations and candidate explanations in physics. European Journal for Philosophy of Science, 10(1):??, January 2020. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-019-0273-5.

Kitcher:2019:LWY

[Kit19] Philip Kitcher. “So … who is your audience?”. European Journal for Philosophy of Science, 9(1):??, January 2019. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-018-0227-3.

Kubiak:2021:ECP

[KK21] Adam P. Kubiak and Paweł Kawalec. The epistemic consequences of pragmatic value-laden scientific inference. European Journal for Philosophy of Science, 11(2):??, June 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-021-00370-9.

Knuuttila:2016:MTW

[KL16] Tarja Knuuttila and Andrea Loettgers. Model templates within and between disciplines: from magnets to gases — and socioeconomic systems. European Journal for Philosophy of Science, 6(3):377–400, October 2016. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-016-0145-1. See cor-

rection [KL23].

Knuuttila:2023:CMT

[KL23] Tarja Knuuttila and Andrea Loettgers. Correction to: Model templates within and between disciplines: from magnets to gases — and socio-economic systems. European Journal for Philosophy of Science, 13(1):??, March 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-022-00508-3. See [KL16].

Kleinhans:2021:EHP

[Kle21] Maarten G. Kleinhans. Down to Earth: History and philosophy of geoscience in practice for undergraduate education. European Journal for Philosophy of Science, 11(3):??, September 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-021-00402-4.

Koskinen:2016:EAT

[KM16a] Inkeri Koskinen and Uskali Mäki. Extra-academic transdisciplinarity and scientific pluralism: what might they learn from one another? European Journal for Philosophy of Science, 6(3):419– 444, October 2016. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-016-0141-5.

Kuorikoski:2016:TAL

[KM16b] Jaakko Kuorikoski and Caterina Marchionni. Triangulation across the lab, the scanner and the field: the case of social preferences. European Journal for Philosophy of Science, 6(3):361–376, October 2016. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-016-0154-0.

Knuuttila:2021:EAM

[Knu21] Tarja Knuuttila. Epistemic artifacts and the modal dimension of modeling. European Journal for Philosophy of Science, 11(3):??, September 2021. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-021-00374-5.

Koberinski:2023:GFS

[Kob23] Adam Koberinski. Generalized frameworks: Structuring searches for new physics. European Journal for Philosophy of Science, 13 (1):??, March 2023. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-022-00504-7.

Koiv:2023:GCT

[Kõi23] Riin Kõiv. Genetically caused trait is an interactive kind. European Journal for Philosophy of Science, 13(3):??, September 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-023-00527-8.

Koons:2021:POQ

[Koo21] Robert C. Koons. Powers ontology and the quantum revolution. European Journal for Philosophy of Science, 11(1):??, March 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-020-00314-9.

Kornmesser:2018:FCP

[Kor18] Stephan Kornmesser. Frames and concepts in the philosophy of science. European Journal for Philosophy of Science, 8(2):225– 251, May 2018. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-017-0183-3.

Koskinen:2017:SBS

[Kos17] Rami Koskinen. Synthetic biology and the search for alternative genetic systems: Taking how-possibly models seriously. European Journal for Philosophy of Science, 7(3):493–506, October 2017. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-017-0176-2.

Koskinen:2019:MRD

[Kos19] Rami Koskinen. Multiple realizability as a design heuristic in biological engineering. European Journal for Philosophy of Science, 9 (1):??, January 2019. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-018-0243-3.

Koskinen:2022:RTE

[Kos22] Inkeri Koskinen. Reactivity as a tool in emancipatory activist research. European Journal for Philosophy of Science, 12(4):??, December 2022. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-022-00487-5.

Kozlov:2023:SEB

[Koz23] Anatolii Kozlov. Scientific experiments beyond surprise and beauty. European Journal for Philosophy of Science, 13(3):??, September 2023. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-023-00536-7.

Katzav:2015:IAC

[KP15] Joel Katzav and Wendy S. Parker. Introduction to Assessing climate models: knowledge, values and policy. European Journal for Philosophy of Science, 5(2):141–148, May 2015. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-015-0117-x.

Krickel:2020:RCP

[Kri20] Beate Krickel. Reply to Cartwright, Pemberton, Wieten: “Mechanisms, laws and explanation”. European Journal for Philosophy of Science, 10(3):??, October 2020. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-020-00312-x. See cor-

rection [Kri21] and [CPW20].

Krickel:2021:CRC

[Kri21] Beate Krickel. Correction to: Reply to Cartwright, Pemberton, Wieten: “Mechanisms, laws and explanation”. European Journal for Philosophy of Science, 11(3):??, September 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-021-00396-z. See [CPW20, Kri20].

Kroon:2011:TDW

[Kro11] Frederick Kroon. Theory-dependence, warranted reference, and the epistemic dimensions of realism. European Journal for Philosophy of Science, 1(2):173–191, May 2011. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-010-0004-4.

Kronfeldner:2021:FWM

[Kro21] Maria Kronfeldner. The freedom we mean: a causal independence account of creativity and academic freedom. European Journal for Philosophy of Science, 11(2):??, June 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-021-00373-6.

Kuorikoski:2012:CSE

[Kuo12] Jaakko Kuorikoski. Contrastive statistical explanation and causal heterogeneity. European Journal for Philosophy of Science, 2 (3):435–452, October 2012. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-012-0050-1.

Laetz:2011:DBS

[Lae11] Brian Laetz. Does the Bayesian solution to the paradox of confirmation really support Bayesianism? European Journal for Philosophy of Science, 1(1):39–46, January 2011. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-010-0007-1.

Lalumera:2023:MMC

[Lal23] Elisabetta Lalumera. Moralisation of medicines: The case of hydroxychloroquine. European Journal for Philosophy of Science, 13(3):??, September 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-023-00542-9.

Lam:2015:POQ

[Lam15] Vincent Lam. Primitive ontology and quantum field theory. European Journal for Philosophy of Science, 5(3):387–397, October 2015. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-015-0113-1.

Lam:2021:CMS

[Lam21] Vincent Lam. Climate modelling and structural stability. European Journal for Philosophy of Science, 11(4):??, December 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-021-00414-0.

Lange:2019:PVW

[Lan19] Marc Lange. The principle of virtual work, counterfactuals, and the avoidance of physics. European Journal for Philosophy of Science, 9(3):??, October 2019. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-019-0256-6.

Landes:2020:VEE

[Lan20] Jürgen Landes. Variety of evidence and the elimination of hypotheses. European Journal for Philosophy of Science, 10(2):??, May 2020. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-019-0272-6.

Laraudogoitia:2022:UMV

[Lar22] Jon Pérez Laraudogoitia. Unmoved movers: a very simple and novel form of indeterminism. European Journal for Philosophy of Science, 12(3):??, September 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-022-00475-9.

Laudisa:2014:ANG

[Lau14] Federico Laudisa. Against the ‘no-go’ philosophy of quantum mechanics. European Journal for Philosophy of Science, 4(1):1–17, January 2014. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-013-0071-4.

Laudisa:2018:SMS

[Lau18] Federico Laudisa. Stop making sense of Bell’s theorem and nonlocality? European Journal for Philosophy of Science, 8(2):293–306, May 2018. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-017-0187-z.

Laudisa:2023:ITV

[Lau23] Federico Laudisa. The information-theoretic view of quantum mechanics and the measurement problem(s). European Journal for Philosophy of Science, 13(2):??, June 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-023-00521-0.

Lazarovici:2018:AF

[Laz18] Dustin Lazarovici. Against fields. European Journal for Philosophy of Science, 8(2):145–170, May 2018. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-017-0179-z.

Lohse:2021:FSM

[LC21] Simon Lohse and Stefano Canali. Follow *the* science? On the marginal role of the social sciences in the COVID-19 pandemic. European Journal for Philosophy of Science, 11(4):??, December 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-021-00416-y.

Lee:2021:MS

[LD21] Jonny Lee and Joe Dewhurst. The mechanistic stance. European Journal for Philosophy of Science, 11(1):??, March 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-020-00341-6.

LeBihan:2012:DSV

[Le 12] Soazig Le Bihan. Defending the Semantic View: what it takes. European Journal for Philosophy of Science, 2(3):249–274, October 2012. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-011-0026-6.

LeBihan:2020:STL

[Le 20] Baptiste Le Bihan. String theory, loop quantum gravity and eternalism. European Journal for Philosophy of Science, 10(2):??, May 2020. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-020-0275-3.

Lusk:2022:NEV

[LE22] Greg Lusk and Kevin C. Elliott. Non-epistemic values and scientific assessment: an adequacy-for-purpose view. European Journal for Philosophy of Science, 12(2):??, June 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-022-00458-w.

Leonelli:2019:WDD

[Leo19] Sabina Leonelli. What distinguishes data from models? European Journal for Philosophy of Science, 9(2):??, May 2019. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-018-0246-0.

Lewens:2019:DAL

[Lew19] Tim Lewens. The division of advisory labour: the case of ‘mitochondrial donation’. European Journal for Philosophy of Science, 9 (1):??, January 2019. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-018-0235-3.

Lombardi:2016:DDV

[LFL16] Olimpia Lombardi, Sebastian Fortin, and Cristian López. Deflating the deflationary view of information. European Journal for Philosophy of Science, 6(2):209–230, May 2016. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-015-0128-7.

Linnemann:2020:MNL

[Lin20] Niels Linnemann. On metaphysically necessary laws from physics. European Journal for Philosophy of Science, 10(2):??, May 2020. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-020-00281-1.

Linford:2021:NLR

[Lin21] Daniel Linford. Neo-Lorentzian relativity and the beginning of the universe. European Journal for Philosophy of Science, 11(4): ??, December 2021. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-021-00417-x.

Lin:2022:KTT

[Lin22] Chia-Hua Lin. Knowledge transfer, templates, and the spillovers. European Journal for Philosophy of Science, 12(1):??, March 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-021-00426-w.

Lisciandra:2017:RAT

[Lis17] Chiara Lisciandra. Robustness analysis and tractability in modeling. European Journal for Philosophy of Science, 7(1):79–95, January 2017. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-016-0146-0.

Losada:2018:HQM

[LL18] Marcelo Losada and Olimpia Lombardi. Histories in quantum mechanics: distinguishing between formalism and interpretation. European Journal for Philosophy of Science, 8(3):367–394, October 2018. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-017-0197-x.

Ladyman:2013:WCS

[LLW13] James Ladyman, James Lambert, and Karoline Wiesner. What is a complex system? European Journal for Philosophy of Science, 3(1):33–67, January 2013. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-012-0056-8.

Landes:2018:ECI

[LOP18] Jürgen Landes, Barbara Osimani, and Roland Poellinger. Epistemology of causal inference in pharmacology. European Journal for Philosophy of Science, 8(1):3–49, January 2018. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-017-0169-1.

Lopez:2019:RPH

[Lóp19] Cristian López. Roads to the past: how to go and not to go backward in time in quantum theories. European Journal for Philosophy of Science, 9(2):??, May 2019. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-019-0250-z.

Lopez:2021:TFT

[Lop21] Cristian Lopez. Three facets of time-reversal symmetry. European Journal for Philosophy of Science, 11(2):??, June 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-021-00355-8.

Lopez:2023:SPS

[Lóp23] Cristian López. Should physical symmetries guide metaphysics? Two reasons why they should maybe not. European Journal for Philosophy of Science, 13(2):??, June 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-023-00525-w.

Lopez:2024:QOT

[Lop24] Cristian Lopez. Quantum ontology without textbooks. Nor overlapping. European Journal for Philosophy of Science, 14(1):??, March 2024. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-024-00573-w.

Lopez-Rubio:2020:BDR

[LR20] Ezequiel López-Rubio. The Big Data razor. European Journal for Philosophy of Science, 10(2):??, May 2020. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-020-00288-8.

Linnemann:2022:MPR

[LR22] Niels Linnemann and James Read. Miracles persist: a reply to Sus. European Journal for Philosophy of Science, 12(1):??, March 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-021-00440-y. See [Sus21].

Lam:2024:APB

[LR24] Vincent Lam and Yannick Rousselot. Anthropocene, planetary boundaries and tipping points: interdisciplinarity and values in Earth system science. European Journal for Philosophy of Science, 14(2):??, June 2024. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-024-00579-4.

Lujan:2021:EBM

[LT21] José Luis Luján and Oliver Todt. Evidence based methodology: a naturalistic analysis of epistemic policies in regulatory science. European Journal for Philosophy of Science, 11(1):??, March 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-020-00340-7.

Luchetti:2022:QIN

[Luc22] Michele Luchetti. The quantification of intelligence in nineteenthcentury craniology: an epistemology of measurement perspective. European Journal for Philosophy of Science, 12(4):??, December 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-022-00485-7.

Luc:2023:MVI

[Luc23] Joanna Luc. Motivationalism vs. interpretationalism about symmetries: some options overlooked in the debate about the relationship between symmetries and physical equivalence. European Journal for Philosophy of Science, 13(3):??, September 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-023-00539-4.

Ludwig:2023:NMR

[Lud23] David Ludwig. From naturalness to materiality: reimagining philosophy of scientific classification. European Journal for Philosophy of Science, 13(1):??, March 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-023-00509-w.

Lukits:2021:SPB

[Luk21] Stefan Lukits. Symmetry and partial belief geometry. European Journal for Philosophy of Science, 11(3):??, September 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-021-00398-x.

Lusk:2022:HVC

[Lus22] Greg Lusk. Is HPS a valuable component of a STEM education? An empirical study of student interest in HPS courses within an undergraduate science curriculum. European Journal for Philosophy of Science, 12(1):??, March 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-021-00433-x.

Lutz:2012:ALP

[Lut12] Sebastian Lutz. Artificial language philosophy of science. European Journal for Philosophy of Science, 2(2):181–203, May 2012. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-011-0042-6.

Landes:2021:PPA

[LWW21] Jürgen Landes, Christian Wallmann, and Jon Williamson. The Principal Principle, admissibility, and normal informal standards of what is reasonable. European Journal for Philosophy of Science, 11(2):??, June 2021. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-021-00352-x.

Machery:2021:MCD

[Mac21a] Edouard Machery. A mistaken confidence in data. European Journal for Philosophy of Science, 11(2):??, June 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-021-00354-9.

MacLeod:2021:AMC

[Mac21b] Miles MacLeod. The applicability of mathematics in computational systems biology and its experimental relations. European Journal for Philosophy of Science, 11(3):??, September 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-021-00403-3.

MacLeod:2021:IPS

[Mac21c] Miles MacLeod. Integrating philosophy of science in civil engineering: an integrative course design strategy. European Journal for Philosophy of Science, 11(4):??, December 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-021-00422-0.

Maki:2016:PIW

[Mäk16] Uskali Mäki. Philosophy of interdisciplinarity. What? Why? How? European Journal for Philosophy of Science, 6(3):327–342, October 2016. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-016-0162-0.

Marchionni:2013:PNH

[Mar13] Caterina Marchionni. Playing with networks: how economists explain. European Journal for Philosophy of Science, 3(3):331– 352, October 2013. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-013-0070-5.

Martin:2019:BLF

[Mar19] Eric C. Martin. “The battle is on”: Lakatos, Feyerabend, and the student protests. European Journal for Philosophy of Science, 9 (2):??, May 2019. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-019-0251-y. See correction [Mar20].

Martin:2020:CBL

[Mar20] Eric C. Martin. Correction to: “The Battle is on”: Lakatos, Feyerabend, and the student protests. European Journal for Philosophy of Science, 10(2):??, May 2020. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-019-0262-8. See [Mar19].

Massimi:2021:CBD

[Mas21a] Michela Massimi. Cosmic Bayes. Datasets and priors in the hunt for dark energy. European Journal for Philosophy of Science, 11 (1):??, March 2021. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-020-00338-1.

Massimi:2021:ITC

[Mas21b] Michela Massimi. Introduction to topical collection on perspectivism in science: metaphysical and epistemological reflections. European Journal for Philosophy of Science, 11(2):??, June 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-021-00359-4.

Mattig:2022:CEE

[Mät22] Peter Mättig. Classifying exploratory experimentation — three case studies of exploratory experimentation at the LHC. European Journal for Philosophy of Science, 12(4):??, December 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-022-00496-4.

Mazzola:2014:CDT

[Maz14] Claudio Mazzola. Can discrete time make continuous space look discrete? European Journal for Philosophy of Science, 4(1):19– 30, January 2014. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-013-0072-3.

Mazzola:2016:SFB

[Maz16] Claudio Mazzola. Still foes: Benovsky on relationism and substantivalism. European Journal for Philosophy of Science, 6(2): 247–260, May 2016. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-015-0132-y.

McCoy:2020:IAB

[McC20] C. D. McCoy. Interpretive analogies between quantum and statistical mechanics. European Journal for Philosophy of Science, 10 (1):??, January 2020. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-019-0268-2.

McLaughlin:2011:PRP

[McL11] Amy McLaughlin. In pursuit of resistance: pragmatic recommendations for doing science within one’s means. European Journal for Philosophy of Science, 1(3):??, October 2011. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-011-0030-x.

Meincke:2019:ABA

[Mei19] Anne Sophie Meincke. Autopoiesis, biological autonomy and the process view of life. European Journal for Philosophy of Science, 9 (1):??, January 2019. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-018-0228-2.

Mennes:2020:PMB

[Men20] Julie Mennes. Putting multidisciplinarity (back) on the map. European Journal for Philosophy of Science, 10(2):??, May 2020. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-020-00283-z.

Miller:2014:SVP

[Mil14] Boaz Miller. Science, values, and pragmatic encroachment on knowledge. European Journal for Philosophy of Science, 4(2):253– 270, May 2014. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-014-0087-4.

Mugg:2021:SRC

[MK21] Joshua Mugg and Muhammad Ali Khalidi. Self-reflexive cognitive bias. European Journal for Philosophy of Science, 11(3):??, September 2021. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-021-00404-2.

Maki:2016:IAP

[MM16] Uskali Mäki and Miles MacLeod. Interdisciplinarity in action: philosophy of science perspectives. European Journal for Philosophy of Science, 6(3):323–326, October 2016. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-016-0161-1.

McCoy:2018:SMD

[MM18] C. D. McCoy and Michela Massimi. Simplified models: a different perspective on models as mediators. European Journal for Philosophy of Science, 8(1):99–123, January 2018. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-017-0178-0.

MacLeod:2016:IPS

[MN16] Miles MacLeod and Nancy J. Nersessian. Interdisciplinary problem- solving: emerging modes in integrative systems biology. European Journal for Philosophy of Science, 6(3):401–418, October 2016. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-016-0157-x.

Mohammadian:2022:ADD

[Moh22] Mousa Mohammadian. An Armstrongian defense of dispositional monist accounts of laws of nature. European Journal for Philosophy of Science, 12(3):??, September 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-022-00481-x.

Molinini:2022:DCA

[Mol22] Daniele Molinini. Direct and converse applications: Two sides of the same coin? European Journal for Philosophy of Science, 12 (1):??, March 2022. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-021-00431-z.

Montano:2013:BSN

[Mon13] Ulianov Montano. Beauty in science: a new model of the role of aesthetic evaluations in science. European Journal for Philosophy of Science, 3(2):133–156, May 2013. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-013-0064-3.

Montuschi:2017:USM

[Mon17] Eleonora Montuschi. Using science, making policy: what should we worry about? European Journal for Philosophy of Science, 7(1):57–78, January 2017. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-016-0143-3.

Morganti:2019:OSR

[Mor19] Matteo Morganti. From ontic structural realism to metaphysical coherentism. European Journal for Philosophy of Science, 9(1): ??, January 2019. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-018-0233-5.

Moruzzi:2021:MCA

[Mor21] Caterina Moruzzi. Measuring creativity: an account of natural and artificial creativity. European Journal for Philosophy of Science, 11 (1):??, March 2021. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-020-00313-w.

Morita:2023:FGD

[Mor23] Kohei Morita. A fine-grained distinction of coarse graining. European Journal for Philosophy of Science, 13(1):??, March 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-023-00513-0.

Martini:2017:MSO

[MP17] Carlo Martini and Manuela Fernández Pinto. Modeling the social organization of science. European Journal for Philosophy of Science, 7(2):221–238, May 2017. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-016-0153-1.

Muller:2023:CSI

[MP23] Thomas Müller and Tomasz Placek. On the condition of Setting Independence. European Journal for Philosophy of Science, 13(4): ??, December 2023. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-023-00550-9.

Menon:2023:SSW

[MS23] Tarun Menon and Jacob Stegenga. Sisyphean science: why value freedom is worth pursuing. European Journal for Philosophy of Science, 13(4):??, December 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-023-00552-7.

Mulder:2018:LH

[Mul18] Jesse M. Mulder. The limits of Humeanism. European Journal for Philosophy of Science, 8(3):671–687, October 2018. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-018-0205-9.

Muller:2024:CSL

[Mül24] Basil Müller. Coordination in social learning: expanding the narrative on the evolution of social norms. European Journal for Philosophy of Science, 14(2):??, June 2024. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-024-00582-9.

Mussgnug:2022:PRM

[Mus22] Alexander Martin Mussgnug. The predictive reframing of machine learning applications: good predictions and bad measurements. European Journal for Philosophy of Science, 12(3):??, September 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-022-00484-8.

Masterton:2017:UCS

[MZG17] George Masterton, Frank Zenker, and Peter Gärdenfors. Using conceptual spaces to exhibit conceptual continuity through scientific theory change. European Journal for Philosophy of Science, 7(1):127–150, January 2017. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-016-0149-x.

Marchionni:2024:RHS

[MZG24] Caterina Marchionni, Julie Zahle, and Marion Godman. Reactivity in the human sciences. European Journal for Philosophy of Science, 14(1):??, March 2024. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-024-00571-y.

Nager:2022:EIC

[Näg22] Paul M. Näger. Evidence for interactive common causes. Resuming the Cartwright–Hausman–Woodward debate. European Journal for Philosophy of Science, 12(1):??, March 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-021-00410-4.

Nappo:2021:CES

[Nap21] Francesco Nappo. Close encounters with scientific analogies of the third kind. European Journal for Philosophy of Science, 11(3):??, September 2021. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-021-00391-4.

Newman:2014:EIW

[New14] Mark Newman. EMU and inference: what the explanatory model of scientific understanding ignores. European Journal for Philosophy of Science, 4(1):55–74, January 2014. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-013-0075-0. See reply

[Kha15].

Ney:2023:TAW

[Ney23] Alyssa Ney. Three arguments for wave function realism. European Journal for Philosophy of Science, 13(4):??, December 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-023-00554-5.

Novick:2023:PO

[NH23] Rose Novick and Philipp Haueis. Patchworks and operations. European Journal for Philosophy of Science, 13(1):??, March 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-023-00515-y.

Nabergall:2019:AHP

[NNW19] Lukas Nabergall, Alejandro Navas, and Eric Winsberg. An antidote for hawkmoths: on the prevalence of structural chaos in non-linear modeling. European Journal for Philosophy of Science, 9(2):??, May 2019. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-018-0244-2.

Norton:2011:HSM

[Nor11] John D. Norton. History of science and the material theory of induction: Einstein’s quanta, Mercury’s perihelion. European Journal for Philosophy of Science, 1(1):3–27, January 2011. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-010-0001-7.

Norton:2018:EHB

[Nor18a] John D. Norton. Erratum to: How to build an infinite lottery machine. European Journal for Philosophy of Science, 8(1):97, January 2018. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-017-0175-3. See [Nor18b].

Norton:2018:HBI

[Nor18b] John D. Norton. How to build an infinite lottery machine. European Journal for Philosophy of Science, 8(1):71–95, January 2018. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/

s13194-017-0174-4. See correction [NP18] and erratum [Nor18a].

Northcott:2022:RF

[Nor22] Robert Northcott. Reflexivity and fragility. European Journal for Philosophy of Science, 12(3):??, September 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-022-00474-w.

Norton:2018:CJD

[NP18] John D. Norton and Alexander R. Pruss. Correction to John D. Norton “How to build an infinite lottery machine”. European Journal for Philosophy of Science, 8(1):143–144, January 2018. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-017-0193-1. See [Nor18b].

Okon:2011:DQM

[OC11] Elias Okon and Craig Callender. Does quantum mechanics clash with the equivalence principle — and does it matter? European Journal for Philosophy of Science, 1(1):133–145, January 2011. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-010-0009-z.

Oddan:2024:RQT

[Odd24] Jessica Oddan. Reconstructions of quantum theory: methodology and the role of axiomatization. European Journal for Philosophy of Science, 14(2):??, June 2024. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-024-00581-w.

OLoughlin:2023:DEC

[O’L23] Ryan O’Loughlin. Diagnosing errors in climate model intercomparisons. European Journal for Philosophy of Science, 13(2):??, June 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-023-00522-z.

Oldofredi:2024:UQI

[Old24] Andrea Oldofredi. Unexpected quantum indeterminacy. European Journal for Philosophy of Science, 14(1):??, March 2024. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-024-00574-9.

Oldofredi:2021:DAQ

[OÖ21] Andrea Oldofredi and Hans Christian Öttinger. The dissipative approach to quantum field theory: conceptual foundations and ontological implications. European Journal for Philosophy of Science, 11(1):??, March 2021. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-020-00330-9.

OConnor:2018:SP

[OW18] Cailin O’Connor and James Owen Weatherall. Scientific polarization. European Journal for Philosophy of Science, 8(3):855– 875, October 2018. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-018-0213-9.

Paetkau:2024:IRE

[Pae24] Tyler Paetkau. Inductive risk and epistemically detrimental dissent in policy-relevant science. European Journal for Philosophy of Science, 14(1):??, March 2024. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-023-00565-2.

Pantsar:2024:TPA

[Pan24] Markus Pantsar. Theorem proving in artificial neural networks: new frontiers in mathematical AI. European Journal for Philosophy of Science, 14(1):??, March 2024. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-024-00569-6.

Parker:2019:SAA

[Par19] Matthew W. Parker. Symmetry arguments against regular probability: a reply to recent objections. European Journal for Philosophy of Science, 9(1):??, January 2019. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-018-0229-1.

Parker:2021:WRW

[Par21] Matthew W. Parker. Weintraub’s response to Williamson’s coin flip argument. European Journal for Philosophy of Science, 11(3): ??, September 2021. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-021-00389-y.

Passon:2019:IFD

[Pas19] Oliver Passon. On the interpretation of Feynman diagrams, or, did the LHC experiments observe H → γγ? European Journal for Philosophy of Science, 9(2):??, May 2019. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-018-0245-1.

Pence:2021:WFR

[Pen21] Charles H. Pence. W. F. R. Weldon changes his mind. European Journal for Philosophy of Science, 11(3):??, September 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-021-00384-3.

Peschard:2011:MSM

[Pes11] Isabelle Peschard. Making sense of modeling: beyond representation. European Journal for Philosophy of Science, 1(3):??, October 2011. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-011-0032-8.

Perez-Escobar:2022:PAM

[PES22] José Antonio Pérez-Escobar and Deniz Sarikaya. Purifying applied mathematics and applying pure mathematics: how a late Wittgensteinian perspective sheds light onto the dichotomy. European Journal for Philosophy of Science, 12(1):??, March 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-021-00435-9.

Peterson:2016:DPA

[Pet16] Martin Peterson. Do pragmatic arguments show too much? European Journal for Philosophy of Science, 6(2):165–172, May 2016. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-015-0120-2.

Peterson:2018:MEC

[Pet18] Clayton Peterson. Methodological empiricism and the choice of measurement models in social sciences. European Journal for Philosophy of Science, 8(3):831–854, October 2018. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-018-0212-x.

Peters:2021:HFE

[Pet21] Uwe Peters. Hidden figures: epistemic costs and benefits of detecting (invisible) diversity in science. European Journal for Philosophy of Science, 11(1):??, March 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-021-00349-6.

Pexton:2014:CAM

[Pex14] Mark Pexton. Can asymptotic models be explanatory? European Journal for Philosophy of Science, 4(2):233–252, May 2014. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-014-0084-7.

Pexton:2016:EIH

[Pex16] Mark Pexton. Emergence and interacting hierarchies in shock physics. European Journal for Philosophy of Science, 6(1):91– 122, January 2016. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-015-0126-9.

Perez-Gonzalez:2020:MEC

[PG20] Saúl Pérez-González. Mechanistic explanations and components of social mechanisms. European Journal for Philosophy of Science, 10 (3):??, October 2020. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-020-00300-1.

Pinna:2015:PQG

[Pin15] Simonluca Pinna. The philosophy of quantum gravity — lessons from Nicolai Hartmann. European Journal for Philosophy of Science, 5(3):279–296, October 2015. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-014-0103-8.

Plutynski:2021:HCC

[Plu21] Anya Plutynski. How is cancer complex? European Journal for Philosophy of Science, 11(2):??, June 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-021-00371-8.

Politi:2021:FMS

[Pol21] Vincenzo Politi. Formal models of the scientific community and the value-ladenness of science. European Journal for Philosophy of Science, 11(4):??, December 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-021-00418-w.

Politi:2024:WOL

[Pol24] Vincenzo Politi. Who ought to look towards the horizon? A qualitative study on the collective social responsibility of scientific research. European Journal for Philosophy of Science, 14(2):??, June 2024. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-024-00580-x.

Potochnik:2011:EU

[Pot11] Angela Potochnik. Explanation and understanding. European Journal for Philosophy of Science, 1(1):29–38, January 2011. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-010-0002-6.

Potters:2019:SPM

[Pot19] Jan Potters. Stabilization of phenomenon and meaning. European Journal for Philosophy of Science, 9(2):??, May 2019. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-019-0247-7.

Potters:2020:PEE

[Pot20] Jan Potters. Perspectivism and the epistemology of experimentation. European Journal for Philosophy of Science, 10(2):??, May 2020. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-020-00289-7.

Povich:2020:MCD

[Pov20] Mark Povich. Modality and constitution in distinctively mathematical explanations. European Journal for Philosophy of Science, 10(3):??, October 2020. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-020-00292-y.

Povich:2021:IEI

[Pov21] Mark Povich. Information and explanation: an inconsistent triad and solution. European Journal for Philosophy of Science, 11(2): ??, June 2021. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-021-00368-3.

Powell:2012:CEL

[Pow12] Russell Powell. Convergent evolution and the limits of natural selection. European Journal for Philosophy of Science, 2(3):355– 373, October 2012. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-012-0047-9.

Pero:2016:VMH

[PS16] Francesca Pero and Mauricio Suárez. Varieties of misrepresentation and homomorphism. European Journal for Philosophy of Science, 6(1):71–90, January 2016. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-015-0125-x.

Pence:2018:ENS

[PS18] Charles H. Pence and Daniel G. Swaim. The economy of nature: the structure of evolution in Linnaeus, Darwin, and the modern synthesis. European Journal for Philosophy of Science, 8(3):435– 454, October 2018. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-017-0194-0.

Pulkkinen:2022:VCM

[PUB22] Karoliina Pulkkinen, Sabine Undorf, and Frida A.-M. Bender. Values in climate modelling: testing the practical applicability of the Moral Imagination ideal. European Journal for Philosophy of Science, 12(4):??, December 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-022-00488-4.

Parker:2018:VEH

[PW18] Wendy S. Parker and Eric Winsberg. Values and evidence: how models make a difference. European Journal for Philosophy of Science, 8(1):125–142, January 2018. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-017-0180-6.

Raatikainen:2021:NKT

[Raa21] Panu Raatikainen. Natural kind terms again. European Journal for Philosophy of Science, 11(1):??, March 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-020-00344-3.

Raerinne:2018:AER

[Rae18] Jani Raerinne. Abstraction in ecology: reductionism and holism as complementary heuristics. European Journal for Philosophy of Science, 8(3):395–416, October 2018. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-017-0191-3.

Raz:2018:EKE

[Räz18] Tim Räz. Euler’s Königsberg: the explanatory power of mathematics. European Journal for Philosophy of Science, 8(3):331– 346, October 2018. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-017-0189-x.

Reydon:2022:HIN

[RE22] Thomas A. C. Reydon and Marc Ereshefsky. How to incorporate non-epistemic values into a theory of classification. European Journal for Philosophy of Science, 12(1):??, March 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-021-00438-6.

Reydon:2024:EKM

[RE24] Thomas A. C. Reydon and Marc Ereshefsky. Ethnobiological kinds and material grounding: comments on Ludwig. European Journal for Philosophy of Science, 14(1):??, March 2024. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-024-00568-7.

Reutlinger:2017:DCT

[Reu17] Alexander Reutlinger. Does the counterfactual theory of explanation apply to non-causal explanations in metaphysics? European Journal for Philosophy of Science, 7(2):239–256, May 2017. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-016-0155-z.

Reutlinger:2020:WEW

[Reu20] Alexander Reutlinger. What is epistemically wrong with research affected by sponsorship bias? The evidential account. European Journal for Philosophy of Science, 10(2):??, May 2020. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-020-00280-2.

Reydon:2021:MCP

[Rey21] Thomas A. C. Reydon. Misconceptions, conceptual pluralism, and conceptual toolkits: bringing the philosophy of science to the teaching of evolution. European Journal for Philosophy of Science, 11(2):??, June 2021. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-021-00363-8.

Rochefort-Maranda:2016:SMS

[RM16] Guillaume Rochefort-Maranda. Simplicity and model selection. European Journal for Philosophy of Science, 6(2):261–279, May 2016. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-016-0137-1.

Roche:2012:WCT

[Roc12] William A. Roche. A weaker condition for transitivity in probabilistic support. European Journal for Philosophy of Science, 2(1):111–118, January 2012. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-011-0033-7.

Roche:2016:CIP

[Roc16] William Roche. Confirmation, increase in probability, and partial discrimination: a reply to Zalabardo. European Journal for Philosophy of Science, 6(1):1–7, January 2016. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-015-0115-z.

Roche:2017:CTH

[Roc17] William Roche. A condition for transitivity in high probability. European Journal for Philosophy of Science, 7(3):435–444, October 2017. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-017-0172-6.

Rosenberg:2013:HJF

[Ros13] Alex Rosenberg. How Jerry Fodor slid down the slippery slope to anti-Darwinism, and how we can avoid the same fate. European Journal for Philosophy of Science, 3(1):1–17, January 2013. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-012-0055-9.

Rovelli:2017:MSA

[Rov17] Carlo Rovelli. Michelangelo’s stone: an argument against platonism in mathematics. European Journal for Philosophy of Science, 7(2):285–297, May 2017. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-016-0159-8.

Romero:2014:PMB

[RP14] Gustavo E. Romero and Daniela Pérez. Presentism meets black holes. European Journal for Philosophy of Science, 4(3):293–308, October 2014. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-014-0085-6.

Ratti:2024:SVT

[RR24] Emanuele Ratti and Federica Russo. Science and values: a twoway direction. European Journal for Philosophy of Science, 14(1): ??, March 2024. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-024-00567-8.

Reydon:2019:ESP

[RTT19] Thomas A. C. Reydon, David Teira, and Adam Toon. EPSA17: Selected papers from the biannual conference in Exeter. European Journal for Philosophy of Science, 9(1):??, January 2019. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-018-0232-6.

Ruiz:2018:CDC

[RU18a] María J. Ferreira Ruiz and Jon Umerez. Correction to: “Dealing with the changeable and blurry edges of living things: a modified version of property-cluster kinds”. European Journal for Philosophy of Science, 8(3):519–520, October 2018. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-018-0217-5. See [RU18b].

Ruiz:2018:DCB

[RU18b] María J. Ferreira Ruiz and Jon Umerez. “Dealing with the changeable and blurry edges of living things: a modified version of property-cluster kinds”. European Journal for Philosophy of Science, 8(3):493–518, October 2018. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-018-0210-z. See cor-

rection [RU18a].

Rubin:2020:RSS

[Rub20] Mark Rubin. “Repeated sampling from the same population?” A critique of Neyman and Pearson’s responses to Fisher. European Journal for Philosophy of Science, 10(3):??, October 2020. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-020-00309-6.

Runhardt:2021:RMD

[Run21] Rosa W. Runhardt. Reactivity in measuring depression. European Journal for Philosophy of Science, 11(3):??, September 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-021-00395-0.

Ruyant:2020:PRN

[Ruy20] Quentin Ruyant. Perspectival realism and norms of scientific representation. European Journal for Philosophy of Science, 10(2):??, May 2020. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-020-00285-x.

Russo:2011:GVS

[RW11] Federica Russo and Jon Williamson. Generic versus single-case causality: the case of autopsy. European Journal for Philosophy of Science, 1(1):47–69, January 2011. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-010-0012-4.

Ryan:2024:CTU

[Rya24] Patrick James Ryan. Cosmic topology, underdetermination, and spatial infinity. European Journal for Philosophy of Science, 14 (2):??, June 2024. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-024-00576-7.

Saatsi:2022:ER

[Saa22] Juha Saatsi. (In)effective realism? European Journal for Philosophy of Science, 12(2):??, June 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-021-00441-x.

Sagrafena:2023:OEP

[Sag23] Cristina Sagrafena. The old evidence problem and the inference to the best explanation. European Journal for Philosophy of Science, 13(1):??, March 2023. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-022-00507-4.

Salessi:2017:APC

[Sal17] Sina Salessi. Aporia of power: On the crises, science, and internal dynamics of the mental health field. European Journal for Philosophy of Science, 7(2):175–200, May 2017. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-016-0151-3.

SanPedro:2012:CMR

[San12] Iñaki San Pedro. Causation, measurement relevance and noconspiracy in EPR. European Journal for Philosophy of Science, 2(1):137–156, January 2012. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-011-0037-3. See reply [Wro14].

SanPedro:2014:MIP

[San14] Iñaki San Pedro. Measurement independence, parameter independence and non-locality. European Journal for Philosophy of Science, 4(3):369–374, October 2014. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-014-0091-8.

Sanjuan:2021:CET

[San21] Mariano Sanjuán. Constitutive elements through perspectival lenses. European Journal for Philosophy of Science, 11(1):??, March 2021. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-020-00333-6.

Sarkar:2013:CCI

[Sar13] Sahotra Sarkar. Carnap and the compulsions of interpretation: Reining in the liberalization of empiricism. European Journal for Philosophy of Science, 3(3):353–372, October 2013. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-013-0073-2.

Sartenaer:2015:SVD

[Sar15] Olivier Sartenaer. Synchronic vs. diachronic emergence: a reappraisal. European Journal for Philosophy of Science, 5(1):31–54, January 2015. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-014-0097-2.

Stokes:2015:MAI

[SB15] Dustin Stokes and Vincent Bergeron. Modular architectures and informational encapsulation: a dilemma. European Journal for Philosophy of Science, 5(3):315–338, October 2015. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-015-0107-z.

Schulz:2013:OEM

[Sch13] Armin W. Schulz. Overextension: the extended mind and arguments from evolutionary biology. European Journal for Philosophy of Science, 3(2):241–255, May 2013. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-013-0066-1.

Scholl:2015:IBE

[Sch15] Raphael Scholl. Inference to the best explanation in the catch-22: how much autonomy for Mill’s method of difference? European Journal for Philosophy of Science, 5(1):89–110, January 2015. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-014-0099-0.

Schindler:2017:TFM

[Sch17] Samuel Schindler. Theoretical fertility McMullin-style. European Journal for Philosophy of Science, 7(1):151–173, January 2017. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-016-0150-4.

Schindler:2022:MLM

[Sch22a] Samuel Schindler. Micro-level model explanation and counterfactual constraint. European Journal for Philosophy of Science, 12 (2):??, June 2022. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-022-00465-x.

Schurz:2022:TCG

[Sch22b] Gerhard Schurz. Tacking by conjunction, genuine confirmation and convergence to certainty. European Journal for Philosophy of Science, 12(3):??, September 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-022-00470-0.

Sanchez-Dorado:2020:NWC

[SD20] Julia Sánchez-Dorado. Novel & worthy: creativity as a thick epistemic concept. European Journal for Philosophy of Science, 10(3): ??, October 2020. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-020-00303-y.

Sterner:2021:MCS

[SD21] Beckett Sterner and Christopher DiTeresi. Making coherent senses of success in scientific modeling. European Journal for Philosophy of Science, 11(1):??, March 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-020-00336-3.

Seifert:2020:SEM

[Sei20] Vanessa A. Seifert. The strong emergence of molecular structure. European Journal for Philosophy of Science, 10(3):??, October 2020. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-020-00308-7.

Sengers:2017:PBH

[Sen17] Geurt Sengers. Presentism and black holes. European Journal for Philosophy of Science, 7(1):1–15, January 2017. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-015-0135-8.

Serban:2017:WCP

[Şer17] Maria Şerban. What can polysemy tell us about theories of explanation? European Journal for Philosophy of Science, 7(1): 41–56, January 2017. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-016-0142-4.

Steel:2018:CMD

[SFG^{+}18a] Daniel Steel, Sina Fazelpour, Kinley Gillette, Bianca Crewe, and Michael Burgess. Correction to: Multiple diversity concepts and their ethical–epistemic implications. European Journal for Philosophy of Science, 8(3):781, October 2018. CO-DEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic).

URL https://link.springer.com/article/10.1007/s13194-

018-0214-8. See [SFG^{+}18b].

Steel:2018:MDC

[SFG^{+}18b] Daniel Steel, Sina Fazelpour, Kinley Gillette, Bianca Crewe, and Michael Burgess. Multiple diversity concepts and their ethical–epistemic implications. European Journal for Philosophy of Science, 8(3):761–780, October 2018. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-018-0209-5. See cor-

rection [SFG^{+}18a].

Sartenaer:2021:WPC

[SGH21] Olivier Sartenaer, Alexandre Guay, and Paul Humphreys. What price changing laws of nature? European Journal for Philosophy of Science, 11(1):??, March 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-020-00327-4.

Schickore:2019:IMI

[SH19] Jutta Schickore and Nora Hangel. “It might be this, it should be that… “ uncertainty and doubt in day-to-day research practice. European Journal for Philosophy of Science, 9(2):??, May 2019. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-019-0253-9.

Shackel:2018:INP

[Sha18] Nicholas Shackel. The Infinity from Nothing paradox and the Immovable Object meets the Irresistible Force. European Journal for Philosophy of Science, 8(3):417–433, October 2018. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-017-0192-2.

Shaw:2021:FFF

[Sha21] Jamie Shaw. Feyerabend, funding, and the freedom of science: the case of traditional Chinese medicine. European Journal for Philosophy of Science, 11(2):??, June 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-021-00361-w.

Shackel:2022:NIP

[Sha22] Nicholas Shackel. The Nothing from Infinity paradox versus Plenitudinous Indeterminism. European Journal for Philosophy of Science, 12(2):??, June 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-022-00463-z.

Shenker:2020:IVE

[She20] Orly Shenker. Information vs. entropy vs. probability. European Journal for Philosophy of Science, 10(1):??, January 2020. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-019-0274-4.

Shevlin:2021:RCI

[She21] Henry Shevlin. Rethinking creative intelligence: comparative psychology and the concept of creativity. European Journal for Philosophy of Science, 11(1):??, March 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-020-00323-8.

Shrapnel:2014:QCE

[Shr14] Sally Shrapnel. Quantum causal explanation: or, why birds fly south. European Journal for Philosophy of Science, 4(3):409–423, October 2014. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-014-0094-5.

Simons:2018:JHB

[Sim18] Massimiliano Simons. The Janus head of Bachelard’s phenomenotechnique: from purification to proliferation and back. European Journal for Philosophy of Science, 8(3):689–707, October 2018. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-018-0206-8.

Simpson:2021:CH

[Sim21] William M. R. Simpson. Cosmic hylomorphism. European Journal for Philosophy of Science, 11(1):??, March 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-020-00342-5.

Slowik:2012:SMP

[Slo12] Edward Slowik. On structuralism’s multiple paths through spacetime theories. European Journal for Philosophy of Science, 2(1): 45–66, January 2012. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-011-0020-z.

Serpico:2023:AVI

[SM23] Davide Serpico and Mariusz Maziarz. Averaged versus individualized: pragmatic N-of-1 design as a method to investigate individual treatment response. European Journal for Philosophy of Science, 13(4):??, December 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-023-00559-0.

Sober:2020:TCC

[Sob20] Elliott Sober. A theory of contrastive causal explanation and its implications concerning the explanatoriness of deterministic and probabilistic hypotheses. European Journal for Philosophy of Science, 10(3):??, October 2020. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-020-00299-5.

Smart:2021:AMP

[SOH21] Paul R. Smart, Kieron O’Hara, and Wendy Hall. Applying mechanical philosophy to web science: The case of social machines. European Journal for Philosophy of Science, 11(3):??, September 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-021-00388-z.

Soler:2011:TAE

[Sol11a] Léna Soler. Tacit aspects of experimental practices: analytical tools and epistemological consequences. European Journal for Philosophy of Science, 1(3):??, October 2011. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-011-0039-1.

Solomon:2011:JPE

[Sol11b] Miriam Solomon. Just a paradigm: evidence-based medicine in epistemological context. European Journal for Philosophy of Science, 1(3):??, October 2011. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-011-0034-6.

Sole:2017:SBT

[Sol17] Albert Solé. Surrealistic Bohmian trajectories appraised. European Journal for Philosophy of Science, 7(3):467–492, October 2017. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-017-0170-8.

Spencer:2018:CCI

[Spe18] Emily Spencer. Conceptual clarification and implicit-association tests: psychometric evidence for racist attitudes. European Journal for Philosophy of Science, 8(1):51–70, January 2018. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-017-0173-5.

Spohn:2014:EAC

[Spo14] Wolfgang Spohn. The epistemic account of ceteris paribus conditions. European Journal for Philosophy of Science, 4(3):385– 408, October 2014. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-014-0093-6.

Sprenger:2013:RBP

[Spr13] Jan Sprenger. The role of Bayesian philosophy within Bayesian model selection. European Journal for Philosophy of Science, 3(1):101–114, January 2013. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-012-0059-5.

Sprenger:2016:PNM

[Spr16] Jan Sprenger. The probabilistic no miracles argument. European Journal for Philosophy of Science, 6(2):173–189, May 2016. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-015-0122-0.

Sprenger:2018:OSB

[Spr18] Jan Sprenger. The objectivity of Subjective Bayesianism. European Journal for Philosophy of Science, 8(3):539–558, October 2018. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-018-0200-1.

Scholl:2013:MCS

[SR13] Raphael Scholl and Tim Räz. Modeling causal structures. European Journal for Philosophy of Science, 3(1):115–132, January 2013. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-012-0060-z.

Sikimic:2022:MLS

[SR22] Vlasta Sikimić and Sandro Radovanović. Machine learning in scientific grant review: algorithmically predicting project efficiency in high energy physics. European Journal for Philosophy of Science, 12(3):??, September 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-022-00478-6.

Schmidt:2015:PPC

[SS15] Gavin A. Schmidt and Steven Sherwood. A practical philosophy of complex climate modelling. European Journal for Philosophy of Science, 5(2):149–169, May 2015. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-014-0102-9.

Staley:2012:SSE

[Sta12] Kent W. Staley. Strategies for securing evidence through model criticism. European Journal for Philosophy of Science, 2(1):21– 43, January 2012. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-011-0022-x.

Stojanovic:2021:CPS

[Sto21a] Milutin Stojanovic. Correction to: Philosophy of sustainability experimentation — experimental legacy, normativity and transfer of evidence. European Journal for Philosophy of Science, 11(4): ??, December 2021. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-021-00432-y. See [Sto21b].

Stojanovic:2021:PSE

[Sto21b] Milutin Stojanovic. Philosophy of sustainability experimentation — experimental legacy, normativity and transfer of evidence. European Journal for Philosophy of Science, 11(3):??, September 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/

s13194-021-00383-4. See correction [Sto21a].

Stromland:2021:MOM

[Str21] Eirik Strømland. Making our “meta-hypotheses” clear: heterogeneity and the role of direct replications in science. European Journal for Philosophy of Science, 11(2):??, June 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-021-00348-7.

Suarez:2014:CEP

[Suá14] Mauricio Suárez. A critique of empiricist propensity theories. European Journal for Philosophy of Science, 4(2):215–231, May 2014. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-014-0083-8.

Suarez:2022:CNE

[Suá22] Mauricio Suárez. The complex nexus of evolutionary fitness. European Journal for Philosophy of Science, 12(1):??, March 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-021-00434-w.

Suarez:2023:MEN

[Suá23] Javier Suárez. Masking, extrinsicness, and the nature of dispositions: the role of niche signals in muscle stem cells. European Journal for Philosophy of Science, 13(2):??, June 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-023-00523-y.

Sullivan:2019:UCC

[Sul19] Emily Sullivan. Universality caused: the case of renormalization group explanation. European Journal for Philosophy of Science, 9 (3):??, October 2019. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-019-0260-x.

Sus:2021:RM

[Sus21] Adán Sus. Relativity without miracles. European Journal for Philosophy of Science, 11(1):??, March 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-020-00311-y. See re-

ply [LR22].

Shan:2021:AEP

[SW21] Yafeng Shan and Jon Williamson. Applying Evidential Pluralism to the social sciences. European Journal for Philosophy of Science, 11(4):??, December 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-021-00415-z.

Stuart:2022:UMU

[SW22] Michael T. Stuart and Daniel Wilkenfeld. Understanding metaphorical understanding (literally). European Journal for Philosophy of Science, 12(3):??, September 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-022-00479-5.

Tambolo:2014:PRF

[Tam14] Luca Tambolo. Pliability and resistance: Feyerabendian insights into sophisticated realism. European Journal for Philosophy of Science, 4(2):197–213, May 2014. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-014-0082-9.

Taylor:2023:APC

[Tay23a] Henry Taylor. Attention as a patchwork concept. European Journal for Philosophy of Science, 13(3):??, September 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-023-00538-5.

Taylor:2023:AAU

[Tay23b] Samuel D. Taylor. Afactivism about understanding cognition. European Journal for Philosophy of Science, 13(3):??, September 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-023-00544-7.

Takacs:2021:FSD

[TB21] Peter Takacs and Pierrick Bourrat. Fitness: static or dynamic? European Journal for Philosophy of Science, 11(4):??, December 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-021-00430-0.

Teh:2015:NRW

[Teh15] Nicholas J. Teh. A note on Rovelli’s ‘why gauge?’. European Journal for Philosophy of Science, 5(3):339–348, October 2015. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-015-0109-x.

Turunen:2023:ESS

[THP23] Petri Turunen, Ilmari Hirvonen, and Ilkka Pättiniemi. Epistemological scientism and the scientific meta-method. European Journal for Philosophy of Science, 13(2):??, June 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-023-00526-9.

Toader:2021:CQM

[Toa21] Iulian D. Toader. On the categoricity of quantum mechanics. European Journal for Philosophy of Science, 11(1):??, March 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-020-00328-3.

Trlifajova:2022:ICA

[Trl22] Kateřina Trlifajová. Infinity and continuum in the alternative set theory. European Journal for Philosophy of Science, 12(1): ??, March 2022. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-021-00429-7.

Tunc:2022:WSR

[Tun22] Duygu Uygun Tunç. We should redefine scientific expertise: an extended virtue account. European Journal for Philosophy of Science, 12(4):??, December 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-022-00498-2.

Taylor:2017:CFR

[TV17] Henry Taylor and Peter Vickers. Conceptual fragmentation and the rise of eliminativism. European Journal for Philosophy of Science, 7(1):17–40, January 2017. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-016-0136-2.

teVrugt:2022:UPI

[tV22] Michael te Vrugt. Understanding probability and irreversibility in the Mori–Zwanzig projection operator formalism. European Journal for Philosophy of Science, 12(3):??, September 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-022-00466-w.

Uebel:2011:CRD

[Ueb11] Thomas Uebel. Carnap’s ramseyfications defended. European Journal for Philosophy of Science, 1(1):71–87, January 2011. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-010-0003-5.

Valente:2014:ATE

[Val14] Giovanni Valente. The approach towards equilibrium in Lanford’s theorem. European Journal for Philosophy of Science, 4(3):309– 335, October 2014. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-014-0086-5.

Valente:2016:PTC

[Val16] Mario Bacelar Valente. Proper time and the clock hypothesis in the theory of relativity. European Journal for Philosophy of Science, 6(2):191–207, May 2016. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-015-0124-y.

VanCamp:2014:EUU

[Van14] Wesley Van Camp. Explaining understanding (or understanding explanation). European Journal for Philosophy of Science, 4(1):95–114, January 2014. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-013-0077-y.

Vassallo:2020:DRG

[Vas20] Antonio Vassallo. Dependence relations in general relativity. European Journal for Philosophy of Science, 10(1):??, January 2020. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-019-0265-5.

Veit:2021:PPA

[VB21] Walter Veit and Heather Browning. Perspectival pluralism for animal welfare. European Journal for Philosophy of Science, 11 (1):??, March 2021. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-020-00322-9.

Vassallo:2017:RAM

[VDE17] Antonio Vassallo, Dirk-André Deckert, and Michael Esfeld. Relationalism about mechanics based on a minimalist ontology of matter. European Journal for Philosophy of Science, 7(2):299– 318, May 2017. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-016-0160-2.

vanDongen:2021:ORW

[vDS21] Noah van Dongen and Michał Sikorski. Objectivity for the research worker. European Journal for Philosophy of Science, 11(3): ??, September 2021. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-021-00400-6.

vanEck:2015:MEE

[vE15] Dingmar van Eck. Mechanistic explanation in engineering science. European Journal for Philosophy of Science, 5(3):349–375, October 2015. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-015-0111-3.

Vessonen:2021:RM

[Ves21] Elina Vessonen. Representation in measurement. European Journal for Philosophy of Science, 11(3):??, September 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-021-00365-6.

Vezer:2017:VER

[Vez17] Martin A. Vezér. Variety-of-evidence reasoning about the distant past. European Journal for Philosophy of Science, 7(2):257–265, May 2017. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-016-0156-y.

Vergara-Fernandez:2023:CME

[VFHS23] Melissa Vergara-Fernández, Conrad Heilmann, and Marta Szymanowska. Contextualist model evaluation: models in financial economics and index funds. European Journal for Philosophy of Science, 13(1):??, March 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-022-00506-5.

Visokolskis:2022:AVS

[VG22] Sandra Visokolskis and Héctor Horacio Gerván. Applied versus situated mathematics in ancient Egypt: bridging the gap between theory and practice. European Journal for Philosophy of Science, 12(1):??, March 2022. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-021-00419-9.

Vaesen:2021:NFT

[VH21] Krist Vaesen and Wybo Houkes. A new framework for teaching scientific reasoning to students from application-oriented sciences. European Journal for Philosophy of Science, 11(2):??, June 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-021-00379-0.

Vickers:2012:HMO

[Vic12] Peter Vickers. Historical magic in old quantum theory? European Journal for Philosophy of Science, 2(1):1–19, January 2012. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-010-0010-6.

Virmajoki:2021:LEH

[Vir21] Veli Virmajoki. Local explanation in historiography of science. European Journal for Philosophy of Science, 11(1):??, March 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-020-00318-5.

Virmajoki:2023:PSF

[Vir23] Veli Virmajoki. Pursuitworthiness in the scheme of futures. European Journal for Philosophy of Science, 13(1):??, March 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-023-00510-3.

Verreault-Julien:2019:UDD

[VJ19] Philippe Verreault-Julien. Understanding does not depend on (causal) explanation. European Journal for Philosophy of Science, 9(2):??, May 2019. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-018-0240-6.

Veit:2021:MAS

[VN21] Walter Veit and Milan Ney. Metaphors in arts and science. European Journal for Philosophy of Science, 11(2):??, June 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-021-00351-y.

Verweij:2022:CTC

[VNT22] Marco Verweij, Steven Ney, and Michael Thompson. Cultural theory’s contributions to climate science: reply to Hansson. European Journal for Philosophy of Science, 12(2):??, June 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-022-00464-y. See [Han20].

Vassend:2017:PSS

[VSF17] Olav Vassend, Elliott Sober, and Branden Fitelson. The philosophical significance of Stein’s paradox. European Journal for Philosophy of Science, 7(3):411–433, October 2017. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-016-0168-7.

Wallace:2023:SBO

[Wal23] David Wallace. The sky is blue, and other reasons quantum mechanics is not underdetermined by evidence. European Journal for Philosophy of Science, 13(4):??, December 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-023-00557-2.

Wang:2024:CSE

[Wan24] Yafeng Wang. Criteria of success for engineering accident investigations: a question-centered account. European Journal for Philosophy of Science, 14(2):??, June 2024. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-024-00578-5.

Wayne:2018:EI

[Way18] Andrew Wayne. Explanatory integration. European Journal for Philosophy of Science, 8(3):347–365, October 2018. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-017-0190-4.

Wilson:2022:PMP

[WB22] Joseph Wilson and F. Garrett Boudinot. Proxy measurement in paleoclimatology. European Journal for Philosophy of Science, 12 (1):??, March 2022. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-021-00444-8.

Weber:2022:CCC

[Web22] Marcel Weber. Coherent causal control: a new distinction within causation. European Journal for Philosophy of Science, 12(4):??, December 2022. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-022-00499-1.

Welch:2011:DTC

[Wel11] John R. Welch. Decision theory and cognitive choice. European Journal for Philosophy of Science, 1(2):147–172, May 2011. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-010-0005-3.

Werndl:2011:OEC

[Wer11] Charlotte Werndl. On the observational equivalence of continuoustime deterministic and indeterministic descriptions. European Journal for Philosophy of Science, 1(2):193–225, May 2011. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-010-0011-5.

Wronski:2017:DBN

[WG17] Leszek Wroński and Michał Tomasz Godziszewski. Dutch books and nonclassical probability spaces. European Journal for Philosophy of Science, 7(2):267–284, May 2017. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-016-0158-9.

Witteveen:2023:TPS

[WG23] Joeri Witteveen and Sara Green. Teaching philosophy of science that matters. European Journal for Philosophy of Science, 13(2): ??, June 2023. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-023-00529-6.

Williamson:2018:JPI

[Wil18] Jon Williamson. Justifying the principle of indifference. European Journal for Philosophy of Science, 8(3):559–586, October 2018. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-018-0201-0.

Winstanley:2022:MPS

[Win22] M. A. Winstanley. Modelling the psychological structure of reasoning. European Journal for Philosophy of Science, 12(2):??, June 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-022-00449-x.

Woodward:2022:MIM

[Woo22] James Woodward. Modeling interventions in multi-level causal systems: supervenience, exclusion and underdetermination. European Journal for Philosophy of Science, 12(4):??, December 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-022-00486-6.

Wolf:2023:RBT

[WR23] William J. Wolf and James Read. Respecting boundaries: theoretical equivalence and structure beyond dynamics. European Journal for Philosophy of Science, 13(4):??, December 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://

link.springer.com/article/10.1007/s13194-023-00545-6.

Wright:2012:MEO

[Wri12] Cory D. Wright. Mechanistic explanation without the ontic conception. European Journal for Philosophy of Science, 2(3):375– 394, October 2012. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-012-0048-8.

Wronski:2014:CSP

[Wro14] Leszek Wroński. On a conjecture by San Pedro. European Journal for Philosophy of Science, 4(3):361–368, October 2014. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-014-0089-2. See [San12].

Wallmann:2020:PPS

[WW20] Christian Wallmann and Jon Williamson. The Principal Principle and subjective Bayesianism. European Journal for Philosophy of Science, 10(1):??, January 2020. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.springer.

com/article/10.1007/s13194-019-0266-4.

Yee:2021:EMS

[Yee21] Adrian K. Yee. Econophysics: making sense of a chimera. European Journal for Philosophy of Science, 11(4):??, December 2021. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-021-00413-1.

Yee:2023:MLM

[Yee23] Adrian K. Yee. Machine learning, misinformation, and citizen science. European Journal for Philosophy of Science, 13(4):??, December 2023. CODEN ???? ISSN 1879-4912 (print), 1879-

4920 (electronic). URL https://link.springer.com/article/
10.1007/s13194-023-00558-1.

Young:2022:EPR

[You22] Mark Thomas Young. From epistemology to policy: reorienting philosophy courses for science students. European Journal for Philosophy of Science, 12(2):??, June 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https://link.

springer.com/article/10.1007/s13194-022-00454-0.

Zach:2022:RAI

[Zac22] Martin Zach. Revisiting abstraction and idealization: how not to criticize mechanistic explanation in molecular biology. European Journal for Philosophy of Science, 12(1):??, March 2022. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (electronic). URL https:

//link.springer.com/article/10.1007/s13194-022-00453-1.

Zahle:2023:RGD

[Zah23] Julie Zahle. Reactivity and good data in qualitative data collection. European Journal for Philosophy of Science, 13(1):??, March 2023. CODEN ???? ISSN 1879-4912 (print), 1879-4920 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s13194-023-00514-z.